### v3

# Algebraic Zhou valuations

Abstract.
In this paper, we generalize Zhou valuations, originally defined on complex domains, to the framework of general schemes. We demonstrate that an algebraic version of the Jonsson–Mustaţă conjecture is equivalent to the statement that every Zhou valuation is quasi-monomial. By introducing a mixed version of jumping numbers and Tian functions associated with valuations, we obtain characterizations of a valuation being a Zhou valuation or computing some jumping number using the Tian functions. Furthermore, we establish the correspondence between Zhou valuations in algebraic settings and their counterparts in analytic settings.
Key words and phrases:
Zhou valuation, Jonsson–Mustaţă’s conjecture, graded sequence of ideals, multiplier ideal, jumping number
2020 Mathematics Subject Classification:
14F18, 12J20, 14B05, 32U05, 32U35
1. Introduction
In the present paper, we work on separated regular connected Noetherian excellent schemes over $k\coloneqq\mathbf{Q}$. We extend the definition of Zhou valuations, originally introduced in [4], from the ring of germs of holomorphic functions on $\mathbf{C}^{n}$ to such schemes, while all the settings are completely algebraic. First, we recall some concepts that are closely related to the definition and properties of algebraic Zhou valuations.
1.1. Background
1.1.1. Jumping number and graded sequence of ideals
Let $X$ be a regular algebraic variety defined over a field with characteristic zero, and let $\mathfrak{a}$ and $\mathfrak{q}$ be nonzero ideals on $X$. The jumping number (associated to $\mathfrak{q}$), denoted by $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$, is defined as
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\coloneqq\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})}{\operatorname{ord}_{E}(\mathfrak{a})}=\min\{\lambda\colon\mathfrak{q}\not\subseteq\mathcal{J}(\mathfrak{a}^{\lambda})\},$
where $E$ runs through all prime divisors over $X$, $\mathcal{J}(\mathfrak{a})$ is the multiplier ideal associated to $\mathfrak{a}$ (cf. [36, Part Three]) and the number $A(\operatorname{ord}_{E})-1$ is the coefficient of the divisor $E$ on $Y$ in the relative canonical class $K_{Y/X}$. When $q=\mathcal{O}_{X}$, the jumping number specializes to the log canonical threshold (see [34, Section 8]), denoted by $\mathrm{lct}(\mathfrak{a})$. Both jumping numbers and log canonical thresholds are important birational invariants in algebraic geometry, which are algebro-geometric analogues of jumping numbers (cf. [19]) and complex singularity exponents (cf. [44, 34, 16]) in complex analysis.
Given a nonzero graded sequence of ideals $\mathfrak{a}_{\bullet}$, one can define an asymptotic jumping number $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$ as
$0<\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\coloneqq\sup_{m}m\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{m})=\lim_{m\to\infty}m\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{m}).$
In [31] Jonsson–Mustaţă showed that if we set $v(\mathfrak{a}_{\bullet})\coloneqq\lim_{m\to\infty}v(\mathfrak{a}_{m})/m$, then
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\inf_{v}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})},$
where $v$ runs all the nontrivial valuations on $X$ and $A(v)$ is the log discrepancy of $v$ (see [31, Section 5] or Section 2.3). Using the compactness arguments, they proved that there must exist a valuation $v$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$ if $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\neq\infty$, i.e. $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\big(A(v)+v(\mathfrak{q})\big)/v(\mathfrak{a}_{\bullet})$.
1.1.2. Jonsson–Mustaţă’s conjectures and Demailly’s strong openness conjecture
In [31], Jonsson–Mustaţă proposed several conjectures regarding the valuations computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. We present here a conjecture closely related to our work; for further versions, see [31, 32].
Conjecture 1.1 (Algebraic version of Jonsson–Mustaţă’s conjecture).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals on $X$ and $\mathfrak{q}$ a nonzero ideal on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<\infty$. There exists a quasi-monomial valuation $v\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$.
While 1.1 was being formulated, the cases for $\dim X\leq 2$ were already established in [31]. Moreover, Xu [46] made a breakthrough on 1.1, that is, proving the case $\mathfrak{q}=\mathcal{O}_{X}$ of 1.1.
One very important purpose of Jonsson–Mustaţă in proposing these conjectures was to study the openness conjecture and the strong openness conjecture through algebraic approaches. In particular, as stated in [32, Theorem D and D’], the validity of 1.1 would imply the strong openness conjecture, and when $\mathfrak{q}=\mathcal{O}_{X}$, the 1.1 regarding the log canonical threshold would imply the openness conjecture.
We recall Demailly–Kollár’s openness conjecture and Demailly’s strong openness conjecture. Let $\varphi$ be a plurisubharmonic (psh for short) function on a complex manifold $M$, and let $\mathcal{J}(\varphi)$ be the multiplier ideal sheaf of $\varphi$ ([38, 39]); see (11.1) for the definition. Set $\mathcal{J}_{+}(\varphi)\coloneqq\bigcup_{\varepsilon>0}\mathcal{J}\big((1+\varepsilon)\varphi\big)$. Then Demailly–Kollár’s openness conjecture ([16]) can be formulated as:
Openness Conjecture
If $\mathcal{J}(\varphi)=\mathcal{O}_{M}$, then $\mathcal{J}_{+}(\varphi)=\mathcal{J}(\varphi)$.
When the restriction of $\mathcal{J}(\varphi)$ is removed, this is Demailly’s strong openness conjecture (see [13, 14]):
Strong Openness Conjecture
$\mathcal{J}_{+}(\varphi)=\mathcal{J}(\varphi)$.
Breakthroughs in the openness conjecture and strong openness conjecture initially emerged from algebraic pathways. Using the valuative tree theory (cf. [20]), Favre–Jonsson [21] (see also [22]) proved the $2$-dimensional case of the openness conjecture, and Jonsson–Mustaţă [31] established the $2$-dimensional case of the strong openness conjecture by proving the $2$-dimensional case of 1.1. Berndtsson [8] proved the openness conjecture by using the complex Brunn-Minkowski inequality (see also [9]). After that, Guan-Zhou [26] proved the strong openness conjecture by movably using the Ohsawa-Takegoshi $L^{2}$ extension theorem ([40]). Meanwhile, Xu’s breakthrough ([46]) on 1.1 for the log canonical threshold case provided an algebraic proof of the openness conjecture without dimensional restrictions.
More precisely, the (algebraic version of) Jonsson-Mustaţă’s conjecture (1.1) can imply the strong openness conjeture through implying the following conjecture ([32, Conjecture B’]):
(Analytic version of) Jonsson-Mustaţă’s conjecture
If $c_{o}^{\mathfrak{q}}(\varphi)<+\infty$, then
$\frac{1}{r^{2}}\mu\big(\{c_{o}^{\mathfrak{q}}(\varphi)\varphi-\log|\mathfrak{q}|<\log r\}\big)$
has a positive lower bound independent of $r\in(0,1]$, where $\varphi$ is a psh germ at $o\in\mathbf{C}^{n}$, $\mathfrak{q}\subseteq\mathcal{O}_{o}$ is a nonzero ideal, $\mu$ is the Lebesgue measure on $\mathbf{C}^{n}$, and
$c_{o}^{\mathfrak{q}}(\varphi)\coloneqq\sup\big\{c\geq 0\colon|\mathfrak{q}|^{2}e^{-2c\varphi}\text{ is locally $L^{1}$ integrable near }o\big\}$
is the jumping number (see [16]).
This (analytic version of) Jonsson-Mustaţă’s conjecture was proved by Guan–Zhou based on the truth of the strong openness conjecture ([27]), and a proof without using the strong openness property can be found in [6], with a sharp effectiveness result related to this conjecture obtained (see also [2]). The effectiveness results of the strong openness conjecture can also be referred to [27] (by giving the $L^{2}$ estimates to the solutions of $\bar{\partial}$-equations), [23] (by the concavity of minimal $L^{2}$ integrals), and [1, 3] (by the log-plurisubharmonicity of fiberwise $\xi$-Bergman kernels, a variant of Berndtsson’s log-plurisubharmonicity of fiberwise Bergman kernels [7]).
Although the strong openness conjecture and the analytic version of Jonsson-Mustaţă’s conjecture have been proved, the algebraic version of Jonsson–Mustaţă’s conjecture still remains open in the case of nontrivial ideal $\mathfrak{q}$ (i.e. $\mathfrak{q}\neq\mathcal{O}_{X}$ in 1.1).
1.1.3. Zhou weight, Zhou number and Zhou valuation
Recall the definitions of Zhou weight, Zhou number and Zhou valuation introduced in [4]. Let $f_{0}=(f_{0,1},\ldots,f_{0,m})$ be a vector of holomorphic functions near the origin $o$ in $\mathbf{C}^{n}$, and let $\varphi_{0}$ be a psh function near $o$ such that $|f_{0}|^{2}e^{-2\varphi_{0}}$ is integrable near $o$.
Definition 1.2 (see [4, Definition 1.2]).
We call that a psh function $\Phi$ near $o$ is a (local) Zhou weight related to $|f_{0}|^{2}e^{-2\varphi_{0}}$ near $o$, if the following three statements hold:
  1. (1)
    $|f_{0}|^{2}e^{-2\varphi_{0}}|z|^{2N_{0}}e^{-2\Phi}$ is integrable near $o$ for large enough $N_{0}\gg 0$;
  2. (2)
    $|f_{0}|^{2}e^{-2\varphi_{0}}e^{-2\Phi}$ is not integrable near $o$;
  3. (3)
    For any psh function $\varphi^{\prime}\geq\Phi+O(1)$ near $o$ such that $|f_{0}|^{2}e^{-2\varphi_{0}}e^{-2\varphi^{\prime}}$ is not integrable near $o$, $\varphi^{\prime}=\Phi+O(1)$ holds.
The relative type of a psh germ $u$ at $o$ to a Zhou weight $\Phi$:
$\sigma(u,\Phi)\coloneqq\sup\{c\geq 0\colon u\leq c\Phi+O(1)\text{ near }o\}$
is called the Zhou number of $u$ to $\Phi$.
Call a valuation $v$ on $\mathcal{O}_{o}$ a Zhou valuation related to $|f_{0}|^{2}e^{-2\varphi_{0}}$, if there exists a Zhou weight $\Phi$ related $|f_{0}|^{2}e^{-2\varphi_{0}}$ near $o$ such that $\sigma(\log|f|,\Phi)=v(f)$ for all $(f,o)\in\mathcal{O}_{o}$.
The existence of Zhou valuations is based on the truth of the strong openness conjecture. By establishing an integral expression of Zhou numbers (see [4, Theorem 1.6]), it was proved in [4] that the Zhou numbers satisfying both tropical additivity and tropical multiplicativity, inducing the valuation structure, which makes that every Zhou weight can be associated with a Zhou valuation (see [4, Corollary 1.9]).
The notions of Zhou weights and Zhou numbers were used to analytically reprove and generalize the characterization of singularities of plurisubharmonic germs established by Boucksom–Favre–Jonsson in [10], while Zhou valuations were also used to characterize if one holomorphic germ can be divided by another (see [4, Theorem 1.11 and Corollary 1.14]). In this context, Tian functions were introduced as technical instruments to facilitate the construction and analysis of these structures. These constructions revealed connections between the singularities in several complex variables and valuation theory.
Furthermore, [5] demonstrated some connections between Zhou valuations and Jonsson–Mustaţă’s conjectures in complex analytic framework. For example, it was showed that for any Zhou valuation $\nu$, there exists a graded sequence of ideals $\mathfrak{a}_{\bullet}$ and a nonzero ideal $\mathfrak{q}$ such that $\nu$ $\mathscr{A}$-computes (see [5, Definition 1.5]) the jumping number $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. The results in [5] reveal not only the possibility but also the necessity of extending the notion of Zhou valuations beyond the analytic setting into a broader algebro-geometric framework.
1.2. Algebraic Zhou valuation and main results
Motivated by these observations, the present paper undertakes a systematic algebraic study of Zhou valuations. The idea of generalizing Zhou valuations originates from the answer in [5] to a question asked by Jonsson, which concerns how to characterize a valuation on $\mathcal{O}_{o}$ appearing as a Zhou valuation (see [5, Question 1.10 and Proposition 1.12]). Based on this characterization, we realize that it is reasonable to define algebraic Zhou valuations related to a given nonzero ideal $\mathfrak{q}$ on more general scheme $X$, as follows:
Definition 1.3 (=Definition 3.2).
A valuation on $X$ is called an algebraic Zhou valuation related to a nonzero ideal $\mathfrak{q}$ if it is a maximal element among the valuations $v$ satisfying $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\leq 1$, where $\mathfrak{a}_{\bullet}^{v}$ is the graded sequence of ideals associated to the valuation $v$ (see (2.1)).
The algebraic Zhou valuations can be seen as the analogue of the original Zhou valuations and we will prove in fact there is direct correspondence in Section 11 (see also Theorem 1.6).
We establish a direct connection between algebraic Zhou valuations and the Jonsson–Mustaţ̦ă conjecture (1.1), and we show that Zhou valuations satisfy the following properties in the most general setting considered in article [31].
Theorem 1.4.
Let $\mathfrak{q}$ be a nonzero ideal on $X$, and $\mathfrak{a}_{\bullet}$ be a graded sequence of ideal on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$. Then the following statements hold.
  1. (1)
    (Corollary 4.2) There exists a Zhou valuation related to $\mathfrak{q}$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$.
  2. (2)
    (Theorem 4.1) Any Zhou valuation $v$ related to $\mathfrak{q}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$, and especially $A(v)=1-v(\mathfrak{q})$.
  3. (3)
    (Theorem 4.6) The 1.1 is equivalent to that all Zhou valuations are quasi-monomial.
  4. (4)
    (Corollary 4.5+Corollary 4.7) If $\mathfrak{q}=\mathcal{O}_{X}$ or $\dim X\leq 2$, then all Zhou valuations related to $\mathfrak{q}$ are quasi-monomial.
By Theorem 1.4, regarding as 1.1, it is natural to investigate the question that how to determine whether a given valuation is a Zhou valuation. To address this question, we develop algebraic analogues of the objects studied in [4], including mixed jumping numbers and (algebraic) Tian functions. For given nonzero ideals $\mathfrak{q}$, $\mathfrak{q}^{\prime}$ and $\mathfrak{a}$, we define the mixed version of jumping number as
$0<\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a})\coloneqq\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+t\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{a})},$
where the infimum is over all divisors $E$ over $X$, and $t\in(-\varepsilon_{0},+\infty)$ with small enough $\varepsilon_{0}>0$ (to allow $t<0$ is very important in the latter discussions). Similarly, for a nonzero graded sequence of ideals $\mathfrak{a}_{\bullet}$, we can then define an asymptotic mixed version jumping number as
$\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\coloneqq\lim_{m\to\infty}m\cdot\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})=\sup_{m}m\cdot\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})\in(0,+\infty].$
For any nonzero ideals $\mathfrak{q}$ and $\mathfrak{q}^{\prime}$ and for any nonzero graded sequence of ideals $\mathfrak{a}_{\bullet}$, the function
$t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$
is called as the (algebraic) Tian function, which is the analogue of the Tian function introduced in [4].
We will establish a criterion, formulated via algebraic Tian functions, for determining whether a given valuation is a Zhou valuation associated to some nonzero ideal. This result can be seen as parallel with [4, Proposition 4.8]. In addition, a characterization on the valuations computing some jumping number will also be demonstrated, while the proof of this result is much farther away from the analytic framework. We are going to prove
Theorem 1.5.
Let $\mathfrak{q}$ be a nonzero ideal on $X$ and $v\in\mathrm{Val}^{*}_{X}$ with $A(v)<+\infty$. Assume $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$. We have the followings:
  1. (1)
    (Proposition 6.3+Proposition 6.5) The valuation $v$ computes the jumping number $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$ if and only if the function
    $t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$
    is linear on $0,+\infty)$ for every nonzero ideal $\mathfrak{q}^{\prime}$ on $X$.
  2. (2)
    (Theorem 8.1+Proposition 8.2) The valuation $v$ is a Zhou valuation related to $\mathfrak{q}$ if and only if $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$ and the function
    $(-\varepsilon_{\mathfrak{q}^{\prime}},+\infty)\ni t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$
    is linear on $0,+\infty)$ and differentiable at $t=0$ for every nonzero ideal $\mathfrak{q}^{\prime}$ on $X$, where $\varepsilon_{\mathfrak{q}^{\prime}}>0$ are sufficiently small.
At last, it is natural to consider if one can go back to the (analytic) Zhou valuations (see Definition 11.4) introduced in [4] from the algebraic Zhou valuations defined in the present paper. In fact, if we let $\mathcal{O}_{o}$ be the local ring of germs of holomorphic functions at the origin $o$ in $\mathbf{C}^{n}$, then restricted to $X=\mathrm{Spec}\,\mathcal{O}_{o}$, the algebraic and analytic Zhou valuations coincide, that is
Theorem 1.6 (=Theorem 11.2).
Let $\mathfrak{q}\subseteq\mathcal{O}_{o}$ be a nonzero ideal, and let $v$ be a valuation on $\mathcal{O}_{o}$ with $v(\mathfrak{m})>0$. Then $v$ is an algebraic Zhou valuation related to $\mathfrak{q}$ if and only if $v$ is an analytic Zhou valuation related to $|\mathfrak{q}|^{2}$.
Actually, Theorem 11.2 can be proved combining with the definition of algebraic Zhou valuation and the results in [4, 5], but for the sake of completeness, we provide a brief proof in Section 11, which involves much more algebraic methods, and is far different with the methods in [4, 5]. We do not assume the existence of Zhou weights in the proof, and the truth of the strong openness property of multiplier ideal sheaves, which is the foundation of the proofs in [4], will only be used in the last step.
To establish the equivalence between these two notions, we start with an (algebraic) Zhou valuation $v$ on $\mathcal{O}_{o}$, and construct the corresponding Zhou weight $\Phi_{v}^{\star}$. Then we show that the Zhou number $\sigma(\log|f|,\Phi_{v}^{\star})$ coincides with $v(f)$. We also use this correspondence to provide algebraic reproofs of several results concerning Zhou weights from [4] (see Corollaries 11.7, 11.8 and 11.9).
Organization
The structure of the paper is as follows. In Section 2, we recall some definitions and properties of valuations and jumping numbers that will be frequently used in subsequent sections. In Section 3, we introduce the Zhou valuations, and in Section 4, we prove Theorem 1.4. To prepare for the proof of Theorem 1.5, we define a mixed version of jumping numbers in Section 5 and extend certain properties of jumping numbers to this setting. Section 6 is devoted to the proof of Theorem 1.5 (1). Section 7 develops technical tools necessary for the proof of Theorem 1.5 (2), via the enlargement of graded sequences, and we prove Theorem 1.5 (2) in Section 8. In Section 9, we show that the cone of Zhou valuations is dense in the space of valuations. In Section 10, we provide characterizations of singularities of graded sequences of ideals via Zhou valuations. Finally, in Section 11, we describe the correspondence between algebraic and analytic Zhou valuations, and complete the proof of Theorem 1.6.
Notations and Conventions
  • We always assume that $X$ is a separated, regular, connected, and Noetherian excellent scheme over $k\coloneqq\mathbf{Q}$. A Noetherian scheme $X$ is excellent if $X$ can be covered by spectra of excellent rings.
  • A log-smooth pair over $X$ is a pair $(Y,D)$ with $Y$ regular and $D$ a reduced effective simple normal crossing (SNC for short) divisor, together with a proper birational morphism $\pi\colon Y\to X$ which is an isomorphism outside $\mathrm{Supp}(D)$. Moreover, we say $(Y,D)\preceq(Y^{\prime},D^{\prime})$ if there is a birational morphism $\phi\colon Y^{\prime}\to Y$ over $X$ such that we have $\mathrm{Supp}(\phi^{*}(D))\subseteq\mathrm{Supp}(D^{\prime})$.
  • Let $\mathfrak{a}\subseteq\mathcal{O}_{X}$ be a nonzero ideal sheaf on $X$. A log resolution of $\mathfrak{a}$ is a proper birational morphism $\pi\colon Y\to X$ such that $Y$ is regular and
    $\pi^{-1}\mathfrak{a}\coloneqq\mathfrak{a}\cdot\mathcal{O}_{Y}=\mathcal{O}_{Y}(-D),$
    where $D$ is an effective divisor on $Y$ such that $D+\mathrm{except}(\pi)$ has simple normal crossing support.
  • All the “Zhou valuation” appears in Section 2 to Section 10 are “algebraic Zhou valuation”, and since both the algebraic Zhou valuation and the analytic Zhou valuation will appear in Section 11, we will make a clear distinction in terminology between the two in this section.
2. Preliminaries
In this section, for the convenience of readers, we recall some basic knowledge about valuations and jumping numbers. Moreover, we have compiled the main results and definitions from [31] that will appear repeatedly in our subsequent sections.
2.1. The space of valuations
2.1.1. Valuations
A real valuation of the function field of $X$ with center $\xi$ on $X$ is a map $v\colon K(X)^{*}\to\mathbf{R}$ satisfying the following:
  1. (1)
    $v(fg)=v(f)+v(g)$ for all $f,g$;
  2. (2)
    $v(f+g)\geq\min\{v(f),v(g)\}$ for all $f,g$;
  3. (3)
    $v(f)=0$ for all $f\in\mathbf{Q}^{*}$;
  4. (4)
    there is an inclusion of local rings $\mathcal{O}_{X,\xi}\hookrightarrow\mathcal{O}_{v}$, where $\mathcal{O}_{X,\xi}$ is the local ring of $\xi\in X$ and $\mathcal{O}_{v}\coloneqq\{f\in K(X)\colon v(f)\geq 0\}$ is the valuation ring associated to the valuation $v$.
We set $v(0)=+\infty$. The trivial valuation is the valuation whose restriction to $K(X)^{*}$ is zero. Note that the center of a valuation, if it exists, is unique as $X$ is separated, and we denote the center of the valuation $v$ by $c_{X}(v)$. Let $\mathrm{Val}_{X,\xi}$ be the set of nontrivial real valuations of $K(X)$ with the center $\xi\in X$. Let $\mathrm{Val}_{X}$ be the set of all real valuations of the function field $K(X)$ of $X$ that admits a center on $X$, and we denote $\mathrm{Val}_{X}^{*}\subseteq\mathrm{Val}_{X}$ the subset of nontrivial valuations with a center on $X$.
For each valuation $v$ with a center $\xi$, if $\mathfrak{a}$ is an ideal sheaf on $X$, then we write
$v(\mathfrak{a})\coloneqq\min_{f\in\mathfrak{a}\cdot\mathcal{O}_{X,\xi}}v(f).$
Then this function satisfies the equations
$v(\mathfrak{a}\cdot\mathfrak{b})=v(\mathfrak{a})+v(\mathfrak{b})\quad\text{and}\quad v(\mathfrak{a}+\mathfrak{b})=\min\{v(\mathfrak{a}),v(\mathfrak{b})\}.$
Moreover, $v(\mathfrak{a})>0$ if and only if $\xi\in V(\mathfrak{a})$. For any $v,w\in\mathrm{Val}_{X}$, we say $v\leq w$ if $v(\mathfrak{a})\leq w(\mathfrak{a})$ for all nonzero ideals on $X$.
2.1.2. Divisorial valuations
We say $E$ is a divisor over $X$, if $E$ is a prime divisor on a normal scheme $Y$ with a proper birational morphism $\pi\colon Y\to X$. For a divisor $E$ over $X$, we define the divisorial valuation $\operatorname{ord}_{E}$ that sends each rational function $f$ in $K(X)^{*}=K(Y)^{*}$ to its order of vanishing along $E$. By resolution of singularities (see [29, 42, 43]), we may always assume that $E$ is a prime divisor on a regular scheme $Y$. A valuation $v\in\mathrm{Val}_{X}$ is a divisorial valuation if there is a divisor $E$ over $X$ and $\lambda\in\mathbf{R}_{>0}$ such that $v(f)=\lambda\cdot\operatorname{ord}_{E}(f)$ for all $f\in K(X)^{*}$.
2.1.3. Quasi-monomial valuations
Let $\pi\colon Y\to X$ be a proper birational morphism with regular and connected $Y$, and let $y=(y_{1},\cdots,y_{r})\in\mathcal{O}_{Y,\eta}^{\oplus r}$ be a system of parameters at a point $\eta\in Y$. For $\alpha,\beta\in\mathbf{R}^{r}$, we set $\langle\alpha,\beta\rangle\coloneqq\sum_{i=1}^{r}\alpha_{i}\beta_{i}$.
Definition 2.1.
Set $Y$ and $y$ as above. The quasi-monomial valuation associated with $y$ and $\alpha=(\alpha_{1},\cdots,\alpha_{r})\in\mathbf{R}^{r}_{\geq 0}$ is a valuation $\mathrm{val}_{\alpha}\colon K(X)\to\mathbf{R}\cup\{+\infty\}$ such that
$\mathrm{val}_{\alpha}(f)=\min\{\langle\alpha,\beta\rangle\colon c_{\beta}\neq 0\}$
for all $f\in\mathcal{O}_{Y,\eta}$, where $f=\sum_{\beta\in\mathbf{Z}^{r}_{\geq 0}}c_{\beta}y^{\beta}$ under the completion of a local ring $\mathcal{O}_{Y,\eta}\to\widehat{\mathcal{O}_{Y,\eta}}$.
Given a quasi-monomial valuation $v=\mathrm{val}_{\alpha}$, the center of $v$ on $Y$ is the generic point $\eta^{\prime}$ of $\bigcap_{i}V(y_{i})$ where $i$ runs through all $i$ such that $\alpha_{i}\neq 0$, hence the center of $v$ on $X$ is $\pi(\eta^{\prime})$; see [31, Proposition 3.1].
Let $v\in\mathrm{Val}_{X}^{*}$ with a center $\xi\in X$. The rational rank of $v$ is defined as
$\mathrm{rat.rk}(v)\coloneqq\dim_{\mathbf{Q}}(v(K(X)^{*})\otimes_{\mathbf{Z}}\mathbf{Q});$
and the transcendence degree of $v$ is defined as
$\mathrm{tr.deg}_{X}(v)\coloneqq\mathrm{tr.deg}(\kappa_{v}/\kappa(\xi)),$
where $\kappa_{v}$ and $\kappa(\xi)$ are the residue fields of the valuation ring $\mathcal{O}_{v}$ and $\mathcal{O}_{X,\xi}$ respectively. Write $\dim(\mathcal{O}_{X,\xi})$ as the Krull dimension of the local ring $\mathcal{O}_{X,\xi}$. The Abhyankar inequality says we have the following inequality for all $v\in\mathrm{Val}_{X,\xi}$ (see [45]):
$\mathrm{rat.rk}(v)+\mathrm{tr.deg}_{X}(v)\leq\dim(\mathcal{O}_{X,\xi}),$
and a valuation satisfying the equality is called an Abhyankar valuation.
Proposition 2.2 (cf.[31, Proposition 3.7, Remark 3.9], [18]).
A valuation $v\in\mathrm{Val}_{X}$ is quasi-monomial if and only if $v$ is Abhyankar. A valuation $v\in\mathrm{Val}_{X}^{*}$ is divisorial if and only if $v$ is quasi-monomial with $\mathrm{rat.rk}(v)=1$.
Given a log-smooth pair $(Y,D)$ over $X$, let $\mathrm{QM}_{\eta}(Y,D)$ be the set of all quasi-monomial valuation $v$ that can be described at the point $\eta\in Y$ with respect to coordinates $y_{1},\cdots,y_{r}$ such that each $y_{i}$ defines at $\eta$ an irreducible component of $D$. We put $\mathrm{QM}(Y,D)\coloneqq\bigcup_{\eta}\mathrm{QM}_{\eta}(Y,D)$ where $\eta$ runs through all generic point of a connected component of the intersection of some of $D_{i}$. For every quasi-monomial valuation $v$, there exists a log-smooth pair $(Y,D)$ over $X$ such that $v\in\mathrm{QM}(Y,D)$ by [31, Remark 3.4].
2.1.4. Topology of the space of valuations
Let $\tau$ be the weakest topology on $\mathrm{Val}_{X}$ such that the evaluation map
$\varphi_{\mathfrak{a}}\colon\mathrm{Val}_{X}\to\mathbf{R}\qquad v\mapsto\varphi_{\mathfrak{a}}(v)\coloneqq v(\mathfrak{a})$
is continuous for all nonzero ideals $\mathfrak{a}$ on $X$. Under this topology, the map
$c_{X}\colon\mathrm{Val}_{X}\to X\qquad\qquad v\mapsto c_{X}(v)$
is anti-continuous, i.e., the inverse image of any open subset is closed. Indeed, if $U\subseteq X$ is an open subset of $X$, let $J$ be the ideal sheaf of $X\setminus U$ with reduced scheme structure. Then
$c_{X}^{-1}(U)=\mathrm{Val}_{U}\coloneqq\{v\in\mathrm{Val}_{X}\colon v(J)=0\},$
because $v(J)\geq 0$ for all real valuation $v$ with a center on $X$ and $v(J)>0$ if and only if $c_{X}(v)\in V(J)=(X\setminus U)_{\mathrm{red}}$.
Applying this topology, we provide more descriptions of $\mathrm{QM}(Y,D)\subseteq\mathrm{Val}_{X}$.
Proposition 2.3 ([31, Section 4.2]).
Let $(Y,D)$ be a log-smooth pair over $X$. Then we have the followings:
  1. (1)
    If $\eta$ is the generic point of a connected component of the intersection of $r$ irreducible components $D_{1},\cdots,D_{r}$ of $D$, then the map
    $\mathrm{QM}_{\eta}(Y,D)\to\mathbf{R}^{r}\qquad v\mapsto(v(D_{1}),\cdots,v(D_{r}))$
    gives a homeomorphism onto the cone $\mathbf{R}^{r}_{\geq 0}$.
  2. (2)
    The space $\mathrm{QM}(Y,D)$ is the union of finitely many closed simplicial cones $\mathrm{QM}_{\eta}(Y,D)$.
2.1.5. Retraction maps and structure theorem
We define the retraction map with respect to a log-smooth pair $(Y,D)$ over $X$ as the continuous map
$r_{Y,D}\colon\mathrm{Val}_{X}\to\mathrm{QM}(Y,D)$
that maps a valuation $v$ to the quasi-monomial valuation $w\coloneqq r_{Y,D}(v)$ satisfying $w(D_{i})=v(D_{i})$ for every irreducible component $D_{i}$ of $D$ by Proposition 2.3 (1).
For every valuation $v\in\mathrm{Val}_{X}$ and any ideal $\mathfrak{a}$ on $X$, we have $r_{Y,D}(v)(\mathfrak{a})\leq v(\mathfrak{a})$ with equality if $(Y,D)$ gives a log resolution of $\mathfrak{a}$ (see [31, Corollary 4.8]). Applying retraction maps, the space of valuations could be achieved as a projective limit of simplicial cone complexes.
Proposition 2.4 (cf. [31, Section 4.4]).
The retraction maps induce a homeomorphism
[…]
such that
  1. (1)
    the set of quasi-monomial valuations is dense in $\mathrm{Val}_{X}$;
  2. (2)
    the set of divisorial valuations is dense in $\mathrm{Val}_{X}$.
2.2. Graded sequences and subadditive systems of ideals
2.2.1. Graded sequences
A graded sequence of ideals $\mathfrak{a}_{\bullet}=(\mathfrak{a}_{m})_{m\in\mathbf{Z}_{>0}}$ is a sequence of ideals on $X$ satisfying $\mathfrak{a}_{p}\cdot\mathfrak{a}_{q}\subseteq\mathfrak{a}_{p+q}$ for all $p,q\geq 1$ ([17]). By convention, we put $\mathfrak{a}_{0}=\mathcal{O}_{X}$, and we always assume that such a sequence is nonzero, i.e., $\mathfrak{a}_{m}\neq 0$ for some $m>0$. Let $\mathfrak{a}_{\bullet}$ be a nonzero graded sequence and let $v\in\mathrm{Val}_{X}$. Then we can define
$v(\mathfrak{a}_{\bullet})\coloneqq\inf_{m}\frac{v(\mathfrak{a}_{m})}{m}=\lim_{\begin{subarray}{c}m\to\infty\\ \mathfrak{a}_{m}\neq 0\end{subarray}}\frac{v(\mathfrak{a}_{m})}{m}\in\mathbf{R}_{\geq 0},$
which is well-defined as we have $v(\mathfrak{a}_{p+q})\leq v(\mathfrak{a}_{p}\cdot\mathfrak{a}_{q})=v(\mathfrak{a}_{p})+v(\mathfrak{a}_{q})$.
For any $v\in\mathrm{Val}_{X}^{*}$ with center $\xi=c_{X}(v)\in X$, we can associate $v$ with a graded sequence $\mathfrak{a}^{v}_{\bullet}=\{\mathfrak{a}^{v}_{m}\}$ of ideals such that for an affine open subset $U\subset X$
$\Gamma(U,\mathfrak{a}_{m}^{v})\coloneqq\left\{\begin{array}[]{ll}\{f\in\Gamma(U,\mathcal{O}_{X})\colon v(f)\geq m\}&\text{if }\xi\in U\\ \Gamma(U,\mathcal{O}_{X})&\text{if }\xi\notin U\end{array}\right$. (2.1)
for any $m\in\mathbf{Z}_{\geq 0}$. Below we present a lemma that is repeatedly used in this article.
Lemma 2.5 ([31, Lemma 2.4]).
Let $\mathfrak{a}_{\bullet}^{v}$ be the graded sequence of ideals associated to a nontrivial valuation $v\in\mathrm{Val}_{X}^{*}$. Then for any $w\in\mathrm{Val}_{X}$, we have
$w(\mathfrak{a}_{\bullet}^{v})=\inf_{\mathfrak{b}}\frac{w(\mathfrak{b})}{v(\mathfrak{b})},$
where $\mathfrak{b}$ ranges over ideals on $X$ for which $v(\mathfrak{b})>0$. In particular, $v(\mathfrak{a}_{\bullet}^{v})=1$, and $w\geq v$ if and only if $w(\mathfrak{a}_{\bullet}^{v})\geq 1$.
2.2.2. Subadditive systems
A subadditive system of ideals $\mathfrak{b}_{\bullet}$ is a one-parameter family $(\mathfrak{b}_{t})_{t\in\mathbf{R}_{>0}}$ of nonzero ideals satisfying $\mathfrak{b}_{s+t}\subseteq\mathfrak{b}_{s}\cdot\mathfrak{b}_{t}$ for all $s,t>0$. Again by convention we put $\mathfrak{b}_{0}=\mathcal{O}_{X}$. Similar to graded sequences, let $\mathfrak{b}_{\bullet}$ be a nonzero graded sequence and let $v\in\mathrm{Val}_{X}$. Then we can define
$v(\mathfrak{b}_{\bullet})\coloneqq\sup_{t}\frac{v(\mathfrak{b}_{t})}{t}=\lim_{t\to\infty}\frac{v(\mathfrak{b}_{t})}{t}\in\mathbf{R}_{\geq 0}\cup\{+\infty\}.$
Given a graded sequence of ideals on $X$, denoted by $\mathfrak{a}_{\bullet}$, the asymptotic multiplier ideals of $\mathfrak{a}_{\bullet}$ are defined as
$\mathfrak{b}_{t}\coloneqq\mathcal{J}(\mathfrak{a}_{\bullet}^{t})\coloneqq\mathcal{J}(\mathfrak{a}_{m}^{t/m}),$
where $m$ is divisible enough and $\mathcal{J}(\mathfrak{a})$ is the multiplier ideal associated to an ideal sheaf $\mathfrak{a}$ (see [17, Definition 1.4] or [36, Definition 9.2.3]).
By the definition of $\mathfrak{b}_{\bullet}\coloneqq(\mathfrak{b}_{t})_{t>0}$ associated to a graded sequence of ideals $\mathfrak{a}_{\bullet}$, $\mathfrak{b}_{\bullet}$ is a subadditive system of ideals and $\mathfrak{a}_{m}\subseteq\mathfrak{b}_{m}$ for all $m\in\mathbf{Z}_{>0}$. Moreover, there are many interesting relations between $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ (e.g. Lemma 2.7), and we will introduce them one by one in the other subsections of this section.
2.3. Log discrepancy and jumping numbers
We now introduce the main objects that we wish to study.
2.3.1. Log discrepancy
The log discrepancy $A(v)$ for every valuation $v\in\mathrm{Val}_{X}$ is defined as follows, by applying the topological structure of $\mathrm{Val}_{X}$ (e.g., Proposition 2.3 and Proposition 2.4).
Let $E$ be a divisor over $X$ with a proper birational morphism $\pi\colon Y\to X$. The log discrepancy associated with the divisorial valuation $\operatorname{ord}_{E}$ is defined as
$A(\operatorname{ord}_{E})\coloneqq\operatorname{ord}_{E}(K_{Y/X})+1,$
where $K_{Y/X}$ is the relative canonical divisor of $\pi$. The log discrepancy $A(\operatorname{ord}_{E})$ depends on the scheme $X$ but does not depend on the choice of $Y$ (cf. [35, pp. 40]). We sometimes denote it by $A_{X}(\operatorname{ord}_{E})$ if there is some ambiguity. Then for a divisorial valuation $v=\lambda\cdot\operatorname{ord}_{E}$ for a rational number $\lambda>0$ and a divisor $E$ over $X$, we define $A(v)\coloneqq\lambda\cdot A(\operatorname{ord}_{E})$.
For every quasi-monomial valuation $v\in\mathrm{QM}(Y,D)$ associated with a log-smooth pair $(Y,D)$, one can define (see [31, (5.1)])
$A_{X}(v)\coloneqq\sum_{i=1}^{N}v(D_{i})\cdot A_{X}(\operatorname{ord}_{D_{i}}),$ (2.2)
where $D_{1},\cdots,D_{N}$ are the irreducible components of $D$. It does not depend on any choices made by [31, Proposition 5.1]. Now for an arbitrary valuation $v\in\mathrm{Val}_{X}$, we set
$A_{X}(v)\coloneqq\sup_{(Y,D)}A_{X}(r_{Y,D}(v))\in\mathbf{R}_{\geq 0}\cup\{\infty\},$
where the supremum is over all log-smooth pairs $(Y,D)$ over $X$. When $v\in\mathrm{Val}_{X}^{*}$, then $A(v)>0$. Moreover, $A(v)$ might be $\infty$ for some $v\in\mathrm{Val}_{X}^{*}$; see [31, Remark 5.12].
Below are some basic properties of the log discrepancy function.
Proposition 2.6 (cf. [31]).
The log discrepancy function $A\colon\mathrm{Val}_{X}\to\mathbf{R}_{\geq 0}\cup\{\infty\}$ for every $v\in\mathrm{Val}_{X}$ is well-defined and satisfies the following:
  1. (1)
    $A(r_{Y,D}(v))\leq A(v)$ for each log-smooth pair $(Y,D)$ over $X$, with equality if and only if $v\in\mathrm{QM}(Y,D)$;
  2. (2)
    if $(Y,D)\preceq(Y^{\prime},D^{\prime})$ for log-smooth pairs over $X$, then for all $v\in\mathrm{Val}_{X}$ we have $A(r_{Y,D}(v))\leq A(r_{Y^{\prime},D^{\prime}}(v))$, with equality if and only if $r_{Y^{\prime},D^{\prime}}\in\mathrm{QM}(Y,D)$;
  3. (3)
    the function $A$ is continuous on each $\mathrm{QM}(Y,D)$ and lower semi-continuous on $\mathrm{Val}_{X}$;
  4. (4)
    if $\mu\colon X^{\prime}\to X$ is a proper birational morphism with $X^{\prime}$ regular, then $A_{X}(v)=A_{X^{\prime}}(v)+v(K_{X^{\prime}/X})$ for all $v\in\mathrm{Val}_{X}$.
Now we can introduce a relation between a graded sequence of ideals $\mathfrak{a}_{\bullet}$ on $X$ and its associated asymptotic multiplier ideals $\mathfrak{b}_{\bullet}$.
Lemma 2.7 (cf. [31, Proposition 2.13, Proposition 6.2]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals and let $\mathfrak{b}_{\bullet}$ be the corresponding system of asymptotic multiplier ideals. Then
  1. (1)
    the system $\mathfrak{b}_{\bullet}$ has controlled growth, that is
    $\frac{\operatorname{ord}_{E}(\mathfrak{b}_{t})}{t}>\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})-\frac{A(\operatorname{ord}_{E})}{t}$
    for every divisor $E$ over $X$ and every $t>0$. Moreover, for every $v\in\mathrm{Val}_{X}$, we have
    $v(\mathfrak{b}_{\bullet})-\frac{A(v)}{t}\leq\frac{v(\mathfrak{b}_{t})}{t}\leq v(\mathfrak{b}_{\bullet}).$
  2. (2)
    $\operatorname{ord}_{E}(\mathfrak{a}_{\bullet})=\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})$ for every divisor $E$ over $X$;
  3. (3)
    $v(\mathfrak{a}_{\bullet})\geq v(\mathfrak{b}_{\bullet})$ for all $v\in\mathrm{Val}_{X}$, with equality if $A(v)<\infty$.
2.3.2. Jumping numbers
For every ideal sheaf $\mathfrak{a}$ and every nonzero ideal sheaf $\mathfrak{q}$ on $X$, the jumping number of $\mathfrak{a}$ with respect to $\mathfrak{q}$ is defined by
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\coloneqq\min\{\lambda\geq 0\colon\mathfrak{q}\not\subseteq\mathcal{J}(\mathfrak{a}^{\lambda})\}.$
By convention, we put $\mathrm{lct}^{\mathfrak{q}}(\mathcal{O}_{X})=\infty$, and when $\mathfrak{q}=\mathcal{O}_{X}$ we write it as $\mathrm{lct}(\mathfrak{a})$. Moreover, the Arnold multiplicity of $\mathfrak{a}$ with respect to $\mathfrak{q}$ is defined as
$\mathrm{Arn}^{\mathfrak{q}}(\mathfrak{a})\coloneqq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})^{-1},$
and again when $\mathfrak{q}=\mathcal{O}_{X}$ we write it as $\mathrm{Arn}(\mathfrak{a})$.
Lemma 2.8 (cf. [31, Lemma 1.7]).
If $\pi\colon Y\to X$ is a log resolution of the nonzero ideal $\mathfrak{a}$, and if $\mathfrak{a}\cdot\mathcal{O}_{Y}=\mathcal{O}_{Y}(-\sum_{i}\alpha_{i}E_{i})$ and $K_{Y/X}=\sum_{i}\kappa_{i}E_{i}$, then for every nonzero ideal $\mathfrak{q}$
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=\min_{\alpha_{i}>0}\frac{A(\operatorname{ord}_{E_{i}})+\operatorname{ord}_{E_{i}}(\mathfrak{q})}{\operatorname{ord}_{E_{i}}(\mathfrak{a})}=\min_{\alpha_{i}>0}\frac{\kappa_{i}+1+\operatorname{ord}_{E_{i}}(\mathfrak{q})}{\alpha_{i}}.$ (2.3)
Moreover, for any ideals $\mathfrak{a}$, $\mathfrak{b}$, $\mathfrak{q}$, $\mathfrak{q}_{1}$ and $\mathfrak{q}_{2}$, we have
  1. (1)
    if $\mathfrak{a}\subseteq\mathfrak{b}$, then $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\leq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b})$;
  2. (2)
    $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{m})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})/m$ for every $m\geq 1$;
  3. (3)
    $\mathrm{lct}^{\mathfrak{q}_{1}+\mathfrak{q}_{2}}(\mathfrak{a})=\min_{i=1,2}\mathrm{lct}^{\mathfrak{q}_{i}}(\mathfrak{a})$.
If the minimum in (2.3) is achieved for $E_{i}$, we say that $\operatorname{ord}_{E_{i}}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$. Moreover, $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$ is a birational invariants, that is, if $\varphi\colon X^{\prime}\to X^{\prime}$ is a proper birational morphism with $X$ and $X^{\prime}$ regular, then we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=\mathrm{lct}^{\mathfrak{q^{\prime}}}(\mathfrak{a^{\prime}}),$
where $\mathfrak{a}^{\prime}\coloneqq\mathfrak{a}\cdot\mathcal{O}_{X^{\prime}}$ and $\mathfrak{q^{\prime}}\coloneqq\mathfrak{q}\cdot\mathcal{O}_{X^{\prime}}(-K_{X^{\prime}/X})$.
For a nonzero graded sequence of ideals $\mathfrak{a}_{\bullet}$ and for a subadditive system of ideals $\mathfrak{b}_{\bullet}$, we can define the jumping numbers of $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ with respect to $\mathfrak{q}$, similar to the definitions of $v(\mathfrak{a}_{\bullet})$ and $v(\mathfrak{b}_{\bullet})$ for any $v\in\mathrm{Val}_{X}$,
$\begin{split}\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\coloneqq\sup_{m}m\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{m})=\lim_{m\to\infty}m\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{m})\in(0,+\infty];\\ \mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})\coloneqq\inf_{t}t\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{t})=\lim_{t\to\infty}t\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{t})\in[0,+\infty).\end{split}$
Furthermore, one can prove the following lemma likely to Lemma 2.8.
Lemma 2.9 (cf. [31]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals and $\mathfrak{b}_{\bullet}$ be a subadditive system of ideals $\mathfrak{b}_{\bullet}$. If $\mathfrak{q}$ is a nonzero ideal, then
$\begin{split}\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})}{\operatorname{ord}_{E}(\mathfrak{a}_{\bullet})};\\ \mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})=\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})},\end{split}$
where $E$ runs through all divisors over $X$ satisfying $\operatorname{ord}_{E}(\mathfrak{a}_{\bullet})\neq 0$ and $\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})\neq 0$, respectively. Moreover, if $\mathfrak{b}_{\bullet}$ is the asymptotic multiplier ideals of $\mathfrak{a}_{\bullet}$, then
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})=\min\{\lambda\geq 0\colon\mathfrak{q}\not\subseteq\mathfrak{b}_{\lambda}\}.$
Here we remind the readers to note that there may not be a divisor $E$ over $X$ that allows the above equations to take the infimum (see [31, Example 8.5]).
We also have the following equations to compute $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$ and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})$.
Lemma 2.10 ([31, Section 6.2]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals and $\mathfrak{b}_{\bullet}$ be a subadditive system of ideals $\mathfrak{b}_{\bullet}$, respectively. If $\mathfrak{q}$ is a nonzero ideal, then
$\begin{split}\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\inf_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})};\\ \mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})=\inf_{\begin{subarray}{c}v\in\mathrm{Val}_{X}^{*}\\ A(v)<\infty\end{subarray}}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{b}_{\bullet})}.\end{split}$
We say that a valuation $v\in\mathrm{Val}_{X}^{*}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$ for a nonzero ideal $\mathfrak{q}$ and a graded sequence of ideals $\mathfrak{a}_{\bullet}$ if $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}$. If $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=+\infty$, then every valuation $v\in\mathrm{Val}_{X}^{*}$ with $A(v)<+\infty$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. For general cases, using the compactness arguments, Jonsson–Mustaţă obtained the existence of valuation computing the jumping numbers:
Theorem 2.11 ([31, Theorem 7.3]).
Let $\{a_{\bullet}\}$ be a graded sequence of ideals on $X$, and $\mathfrak{q}$ be a nonzero ideal. Then there exists a valuation $v\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$.
The following lemma is from [31], but there are some typos in the statements there, so we demonstrate a proof below.
Lemma 2.12 ([31, Proposition 7.10]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals on $X$. Assume that $v\in\mathrm{Val}_{X}^{*}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$. Then
  1. (1)
    $v$ also computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$;
  2. (2)
    $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=v(\mathfrak{a}_{\bullet})\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$;
  3. (3)
    any $v^{\prime}\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$ also computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$.
Proof.
For every $w\in\mathrm{Val}_{X}^{*}$, since $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$, we have
$\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\leq\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a}_{\bullet})}.$ (2.4)
To prove (1), we need to prove that for every $w\in\mathrm{Val}_{X}^{*}$, it holds that
$\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet}^{v})}\leq\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a}_{\bullet}^{v})}.$ (2.5)
We can assume $A(v)+v(\mathfrak{q})>0$. Since $v(\mathfrak{a}_{\bullet}^{v})=1$ and it follows from (2.4) that
$\frac{A(w)+w(\mathfrak{q})}{A(v)+v(\mathfrak{q})}\geq\frac{w(\mathfrak{a}_{\bullet})}{v(\mathfrak{a}_{\bullet})}\geq w(\mathfrak{a}_{\bullet}^{v}).$
Thus, the inequality (2.5) is verified for every $w\in\mathrm{Val}_{X}^{*}$, which implies (1).
By (1), we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}=\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet}^{v})}\cdot\frac{1}{v(\mathfrak{a}_{\bullet})}=\frac{\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})}{v(\mathfrak{a}_{\bullet})},$
which yields (2).
Assume that $v^{\prime}\in\mathrm{Val}_{X}^{*}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$. Then we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=\frac{A(v^{\prime})+v^{\prime}(\mathfrak{q})}{v^{\prime}(\mathfrak{a}_{\bullet}^{v})}.$
The above equality and (2) induce
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\frac{\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})}{v(\mathfrak{a}_{\bullet})}=\frac{A(v^{\prime})+v^{\prime}(\mathfrak{q})}{v^{\prime}(\mathfrak{a}_{\bullet})}\cdot\frac{v^{\prime}(\mathfrak{a}_{\bullet})}{v(\mathfrak{a}_{\bullet})\cdot v^{\prime}(\mathfrak{a}_{\bullet}^{v})}\geq\frac{A(v^{\prime})+v^{\prime}(\mathfrak{q})}{v^{\prime}(\mathfrak{a}_{\bullet})}.$
This indicates that $v^{\prime}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. ∎
2.4. Jonsson–Mustaţă’s conjectures
In this section, we recall the so-called Jonsson–Mustaţă’s conjectures posed in [31] (see also [32]), with some reformulations for simplicity.
Conjecture 2.13 ([31, Conjecture B], [32, Conjecture C]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals on $X$ with $\mathrm{lct}(\mathfrak{a}_{\bullet})<\infty$.
  • Weak version: there exists a quasi-monomial valuation $v\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}(\mathfrak{a}_{\bullet})$;
  • Strong version: any valuation $v\in\mathrm{Val}_{X}^{*}$ computing $\mathrm{lct}(\mathfrak{a}_{\bullet})$ must be quasi-monomial.
Conjecture 2.14 ([31, Conjecture C, Conjecture 7.4], [32, Conjecture C’]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals on $X$ and $\mathfrak{q}$ a nonzero ideal on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<\infty$.
  • Weak version: there exists a quasi-monomial valuation $v\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$;
  • Strong version: any valuation $v\in\mathrm{Val}_{X}^{*}$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$ must be quasi-monomial.
The following theorem shows that, to prove the above Jonsson–Mustaţă’s conjectures, it suffices to consider the graded sequences associated to valuations in $\mathrm{Val}_{X}^{*}$.
Theorem 2.15 ([31, Theorem 7.7, Theorem 7.8]).
If $v\in\mathrm{Val}_{X}^{*}$ is a nontrivial valuation with $A(v)<+\infty$ and $\mathfrak{q}$ is a nonzero ideal on $X$, the following are equivalent:
  1. (1)
    there is a graded sequence of ideals $\mathfrak{a}_{\bullet}$ on $X$ such that $v$ computes the jumping number $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<\infty$;
  2. (2)
    there is a subadditive system of ideals $\mathfrak{b}_{\bullet}$ of controlled growth such that $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})<\infty$;
  3. (3)
    for every $w\in\mathrm{Val}_{X}$ with $w\geq v$, we have $A(w)+w(\mathfrak{q})\geq A(v)+v(\mathfrak{q})$;
  4. (4)
    $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$.
The $2$-dimensional cases of the conjectures are known to be true, and the conjectures are trivial for dimension $1$.
Theorem 2.16 ([31, Section 9]).
The strong versions of 2.13 and 2.14 are valid for all $X$ with $\dim(X)\leq 2$.
In [46], Xu made a breakthrough on the weak version of 2.13.
Theorem 2.17 ([46, Theorem 1.1]).
The weak version of 2.13 holds for all $X$.
3. Definition and existence of Zhou valuations
This section is devoted to the introduction of Zhou valuations and their associated cone, along with several elementary propositions related to them.
Let $\mathfrak{q}$ be a nonzero ideal on $X$, and denote
$\mathrm{Val}(X;\mathfrak{q})\coloneqq\{v\in\mathrm{Val}_{X}\colon\mathfrak{q}\not\subseteq\mathcal{J}(\mathfrak{a}_{\bullet}^{v})\},$ (3.1)
which is a subset of $\mathrm{Val}_{X}$, where $\mathcal{J}(\mathfrak{a}_{\bullet}^{v})$ is the asymptotic multiplier ideal of the graded sequence $\mathfrak{a}^{v}_{\bullet}$. In case that $\mathfrak{q}$ is nonzero, clearly the set $\mathrm{Val}(X;\mathfrak{q})$ is nonempty. By the definition of jumping numbers, one obtains the following description of $\mathrm{Val}(X;\mathfrak{q})$.
Lemma 3.1.
For any nonzero ideal $\mathfrak{q}$ on $X$,
$\mathrm{Val}(X;\mathfrak{q})=\{v\in\mathrm{Val}_{X}\colon\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\leq 1\}.$
Moreover, for every $v\in\mathrm{Val}(X;\mathfrak{q})$, we have $v\leq\operatorname{ord}_{\xi}$, where
$\operatorname{ord}_{\xi}(f)=\sup\big\{r\colon f\in\mathfrak{m}^{r}\cdot\mathcal{O}_{X,\xi}\big\},\quad\forall\,f\in\mathcal{O}_{X,\xi},$
$\xi=c_{X}(v)$, and $\mathfrak{m}$ is the ideal defining $\bar{\xi}$ with the reduced scheme structure.
Proof.
By Lemma 2.9, we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})=\min\{\lambda\geq 0\colon\mathfrak{q}\not\subseteq\mathcal{J}(\lambda\cdot\mathfrak{a}^{v}_{\bullet})\},$ (3.2)
where $\mathcal{J}(\lambda\cdot\mathfrak{a}^{v}_{\bullet})$ is the asymptotic multiplier ideal with coefficient $\lambda$ (see [36]). Thus, it is clear that $\mathrm{Val}(X;\mathfrak{q})$ consists of these $v\in\mathrm{Val}_{X}$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\leq 1$.
If $v\in\mathrm{Val}(X;\mathfrak{q})$, then by Theorem 2.11 we can pick $w\in\mathrm{Val}_{X}^{*}$ which computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})$, that is
$1\geq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})=\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a}^{v}_{\bullet})}.$
After rescaling $w$, we can let $w(\mathfrak{a}^{v}_{\bullet})=1$, which implies $w\geq v$. Then $A(w)\leq 1$. Finally, according to Izumi’s inequality (ref. [30] or [31, Proposition 5.10]), we get
$v\leq w\leq A(w)\operatorname{ord}_{c_{X}(w)}\leq\operatorname{ord}_{c_{X}(w)}\leq\operatorname{ord}_{\xi},$
due to $c_{X}(w)\in\overline{c_{X}(v)}=\bar{\xi}$ by $v\leq w$ (ref. [31, Lemma 4.4]). ∎
Now we give the definition of Zhou valuations on $X$.
Definition 3.2.
A valuation $v\in\mathrm{Val}_{X}$ is called a Zhou valuation related to $\mathfrak{q}$ if and only if $v$ is a maximal element in the set $\mathrm{Val}(X;\mathfrak{q})$, i.e., there is no $w\in\mathrm{Val}(X;\mathfrak{q})$ such that $w\geq v$ and $w\neq v$.
According to Zorn’s lemma, the Zhou valuation related to $\mathfrak{q}$ always exists for nonzero $\mathfrak{q}$ (but it can be not unique). More precisely, we have the following theorem.
Theorem 3.3.
Fix a nonzero ideal $\mathfrak{q}$ on $X$. For any graded sequence $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{m}\}$ satisfying $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=1$, there exists a Zhou valuation $v$ related to $\mathfrak{q}$ such that $\mathfrak{a}_{m}\subseteq\mathfrak{a}_{m}^{v}$ for each $m\in\mathbf{Z}_{+}$.
Proof.
Denote
$\mathcal{S}(\mathfrak{a}_{\bullet})\coloneqq\big\{w\in\mathrm{Val}(X;\mathfrak{q})\colon\mathfrak{a}_{m}\subseteq\mathfrak{a}_{m}^{w},\ \forall\,m\big\}.$
First, we prove $\mathcal{S}(\mathfrak{a}_{\bullet})\neq\emptyset$. By Theorem 2.11, we can pick $w\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$, i.e.,
$1=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a}_{\bullet})},$
and after rescaling $w$, we can let $w(\mathfrak{a}_{\bullet})=1$. This yields $w(\mathfrak{a}_{m})\geq m$ by the definition of $w(\mathfrak{a}_{\bullet})$ and then $\mathfrak{a}_{m}\subseteq\mathfrak{a}_{m}^{w}$ for each $m$. In addition, Lemma 2.12 (2) shows that $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=w(\mathfrak{a}_{\bullet})\cdot\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=1$. Now we get $w\in\mathcal{S}(\mathfrak{a}_{\bullet})$, which shows that the set $\mathcal{S}(\mathfrak{a}_{\bullet})$ is not empty.
Next, by Zorn’s lemma and by noting that for any $w\in\mathrm{Val}(X;\mathfrak{q})$ with $w\geq v$ for some $v\in\mathcal{S}(\mathfrak{a}_{\bullet})$ we must have $w\in\mathcal{S}(\mathfrak{a}_{\bullet})$, to prove the existence of the Zhou valuation, we only need to prove that for any increasing chain in $\mathcal{S}(\mathfrak{a}_{\bullet})$ there is an upper bound belonging to $\mathcal{S}(\mathfrak{a}_{\bullet})$. Suppose $\{v_{j}\}_{j\in\mathbf{Z}_{+}}\subseteq\mathcal{S}(\mathfrak{a}_{\bullet})$ is an increasing chain, i.e. $v_{j}\leq v_{j+1}$ for any $j$. Set
$\tilde{v}(\mathfrak{q}^{\prime})\coloneqq\sup_{j}v_{j}(\mathfrak{q}^{\prime}),\quad\forall\text{ ideal }\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{X}.$
[…]
$\mathfrak{q}\not\subseteq\mathcal{J}\left(\frac{1}{m}\cdot\mathfrak{a}^{\tilde{v}}_{m}\right),\quad\forall\,m\in\mathbf{Z}_{+}.$
Since $\big(\mathfrak{a}_{m}^{v_{j}}\big)_{j\in\mathbf{Z}_{+}}$ is an increasing sequence of ideals on $X$ with respect to $j$, by Noetherian property, there exists some $j_{m}$ such that $\mathfrak{a}^{\tilde{v}}_{m}=\mathfrak{a}^{v_{j_{m}}}_{m}$. Meanwhile, it follows from $v_{j_{m}}\in\mathrm{Val}(X;\mathfrak{q})$ that
[…]
Thus, $\tilde{v}\in\mathrm{Val}(X;\mathfrak{q})$ is an upper bound of the increasing chain $\{v_{j}\}$. Then the proof is done by Zorn’s lemma. ∎
Particularly, when the graded sequence in Theorem 3.3 is associated to a valuation, we immediately get the following corollary.
Corollary 3.4.
For any nonzero ideal $\mathfrak{q}$ on $X$ and any $w\in\mathrm{Val}_{X}$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=1$, there exists a Zhou valuation $v$ related to $\mathfrak{q}$ such that $v\geq w$.
Proof.
Note that $\mathfrak{a}_{m}^{w}\subseteq\mathfrak{a}_{m}^{v}$ for each $m$ induces $v(\mathfrak{a}_{\bullet}^{w})\geq 1$ and then $v\geq w$. Thus, the needed Zhou valuation comes from Theorem 3.3. ∎
It is easy to see
Proposition 3.5.
Every Zhou valuation $v$ related to $\mathfrak{q}$ must satisfy $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$.
Proof.
Suppose $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=c<1$ for a Zhou valuation $v$ related to $\mathfrak{q}$. Then we have $v/c\in\mathrm{Val}(X;\mathfrak{q})$ and $v/c\geq v$. This contradicts to the definition of Zhou valuation. ∎
By multiplying by a positive scalar product, we extend the definition of Zhou valuation to a cone space.
Definition 3.6.
For every nonzero ideal $\mathfrak{q}$ on $X$, we denote by $\mathrm{ZVal}_{X}(\mathfrak{q})$ the set of all Zhou valuations on $X$ related to $\mathfrak{q}$, and denote the cone of Zhou valuations on $X$ by
$\mathrm{ZVal}_{X}\coloneqq\{v=cv^{\prime}\in\mathrm{Val}_{X}\colon c\in\mathbf{R}_{+},\ v^{\prime}\in\mathrm{ZVal}_{X}(\mathfrak{q})\}.$
Then Corollary 3.4 directly implies
Proposition 3.7.
Let $\mathfrak{q}$ be a nonzero ideal on $X$, and $v\in\mathrm{Val}_{X}^{*}$ satisfying $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})<+\infty$. Then there exists $v^{\prime}\in\mathrm{ZVal}_{X}$ such that $v^{\prime}\geq v$ and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v^{\prime}})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$.
Proof.
Note that $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\cdot v\in\mathrm{Val}(X;\mathfrak{q})$. By Corollary 3.4, there exists a Zhou valuation $w$ related to $\mathfrak{q}$ such that $w\geq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\cdot v$. Let $v^{\prime}=w/\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$. Then $v^{\prime}\in\mathrm{ZVal}_{X}$ satisfying $v^{\prime}\geq v$ and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v^{\prime}})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$. ∎
4. Zhou valuations computing jumping numbers
The goal of this section is to establish the proof of Theorem 1.4. First, we compute the log discrepancy of Zhou valuation.
Theorem 4.1.
Let $v$ be a Zhou valuation related to $\mathfrak{q}$. Then the log discrepancy
$A(v)=1-v(\mathfrak{q}).$
Moreover, $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$, and every valuation in $\mathrm{Val}^{*}_{X}$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$ is a positive scalar product of $v$.
Proof.
Pick any $w\in\mathrm{Val}^{*}_{X}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$. Then
$1=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a}_{\bullet}^{v})}.$ (4.1)
After rescaling, we may assume $w(\mathfrak{a}_{\bullet}^{v})=1$. Then Lemma 2.12 implies $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=1$, and thus $w\in\mathrm{Val}(X;\mathfrak{q})$. It follows from $w(\mathfrak{a}_{\bullet}^{v})=1$ that $w\geq v$. Since $v$ is maximal in $\mathrm{Val}(X;\mathfrak{q})$, we get $w=v$. Consequently, $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$, and $A(v)+v(\mathfrak{q})=1$ by (4.1). In addition, from the above we can know that every valuation in $\mathrm{Val}^{*}_{X}$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$ is a positive scalar product of $v$. The proof is done. ∎
By Theorem 4.1, for every $v\in\mathrm{ZVal}_{X}$ we have $A(v)<+\infty$, and we repeatedly use this fact in the paper.
Corollary 4.2.
For every nonzero ideal $\mathfrak{q}$ and graded sequence of ideals $\mathfrak{a}_{\bullet}$ on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$, there exists a Zhou valuation $v$ related to $\mathfrak{q}$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$.
Proof.
Pick any $w\in\mathrm{Val}^{*}_{X}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. We may assume $w(\mathfrak{a}_{\bullet})=1$. By Lemma 2.12, $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$, and $w$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})$. Then $w/\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\in\mathrm{Val}(X;\mathfrak{q})$, and we can take a Zhou valuation related to $\mathfrak{q}$ such that $v\geq w/\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. The above and Theorem 4.1 indicate
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\frac{A(v)+v(\mathfrak{q})}{1/\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})}\geq\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet}^{w})}.$
Thus, $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})$, which is followed by that $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$; see Lemma 2.12 again. ∎
Notably, if $v\in\mathrm{Val}_{X}$ satisfies $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$ and only positive scalars of $v$ compute $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$, then $v$ is a Zhou valuation related to $\mathfrak{q}$. Indeed, if $v$ were not such a valuation, then by Corollary 4.2, there would exist a Zhou valuation $\tilde{v}$ related to $\mathfrak{q}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$. We would then have $\tilde{v}=v$, contradicting our assumption.
Corollary 4.3.
If $\mathfrak{a}_{\bullet}$ is a graded sequence of ideals, then for every nonzero ideal $\mathfrak{q}$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$ we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})=\min_{v\in\mathrm{ZVal}_{X}}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}.$
Proof.
This is a direct consequence of Corollary 4.2 and Lemma 2.10. ∎
We present a simple and special example to motivate the relationship between Zhou valuations and quasi-monomial valuations.
Example 4.4.
Let $\mathfrak{a}$ and $\mathfrak{q}$ be nonzero ideals on $X$, where $\mathfrak{a}\neq\mathcal{O}_{X}$. We may assume $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=1$. Set $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{m}\}$ with $\mathfrak{a}_{m}=\mathfrak{a}^{m}$. Then $v(\mathfrak{a}_{\bullet})=v(\mathfrak{a})$ for each $v\in\mathrm{Val}_{X}$. We demonstrate all the Zhou valuations $w$ related to $\mathfrak{q}$ with $\mathfrak{a}_{m}\subseteq\mathfrak{a}_{m}^{w}$ for every $m$ in the following.
Pick a log-smooth pair $(Y,E)$ over $X$ which gives a log resolution of $\mathfrak{a}\cdot\mathfrak{q}$. It is known (see [31, Lemma 6.7]) that
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=\min_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a})},$
where the equality is achieved for $v$ if and only if $v\in\mathrm{QM}(Y,E)$ and $\operatorname{ord}_{E_{i}}$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$ for every irreducible component $E_{i}$ of $E$ for which $v(E_{i})>0$.
Note that any Zhou valuation $w$ related to $\mathfrak{q}$ with $\mathfrak{a}_{m}\subseteq\mathfrak{a}_{m}^{w}$ for each $m$ (equivalent to $w(\mathfrak{a})\geq 1$) must compute $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$, since (by Theorem 4.1)
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=1=A(w)+w(\mathfrak{q})\geq\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a})}.$
It follows that $w\in\mathrm{QM}(Y,E)$ and satisfies the above conditions. More precisely, if we let $w=\mathrm{val}_{\alpha}\in\mathrm{QM}(Y,E)$, with $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbf{R}_{\geq 0}$ and $\alpha_{i}=\mathrm{val}_{\alpha}(E_{i})$, then for each $i$ with $\alpha_{i}>0$, $\operatorname{ord}_{E_{i}}$ must compute $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$. Also, we have $\mathrm{val}_{\alpha}(\mathfrak{a})=1$ since $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=1$ (Lemma 2.12).
Conversely, we can show that any $\mathrm{val}_{\alpha}\in\mathrm{QM}(Y,E)$ with
$\{i\colon\alpha_{i}\coloneqq\mathrm{val}_{\alpha}(E_{i})\neq 0\}\subseteq\{i\colon\operatorname{ord}_{E_{i}}\text{ computes }\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\}$
and $\mathrm{val}_{\alpha}(\mathfrak{a})=1$ is a Zhou valuation related to $\mathfrak{q}$, i.e. every valuation that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$ is Zhou valuation related to $\mathfrak{q}$ (with some rescaling). Otherwise, let $v^{\prime}\geq\mathrm{val}_{\alpha}$ be a Zhou valuation related to $\mathfrak{q}$. Then $v^{\prime}(\mathfrak{a})\geq v_{\alpha}(\mathfrak{a})=1$. Repeating the above arguments, we can find that $v^{\prime}\in\mathrm{QM}(Y,E)$ and $v^{\prime}=\mathrm{val}_{\alpha}$ eventually.
Noting that in case $\mathfrak{q}=\mathcal{O}_{X}$, we have the following result, which is based Xu’s solution (see Theorem 2.17) to the weak version of Jonsson–Mustaţă’s conjecture.
Corollary 4.5.
Any Zhou valuation related to $\mathcal{O}_{X}$ is quasi-monomial.
Proof.
Let $v$ be a Zhou valuation related to $\mathcal{O}_{X}$. By [46, Theorem 1.1], there exists a quasi-monomial valuation $w\in\mathrm{Val}_{X}^{*}$ computing $\mathrm{lct}(\mathfrak{a}_{\bullet}^{v})$. Thus, by Theorem 4.1, there exists $\lambda>0$ such that $v=\lambda\cdot w$, which implies that $v$ is also quasi-monomial. ∎
For the general cases, we can see that the weak version of Jonsson–Mustaţă’s conjecture is equivalent to that every Zhou valuation is quasi-monomial, where we denote by $\mathrm{QM}_{X}$ the set of quasi-monomial valuations on $X$ in the following.
Theorem 4.6.
The following statements are equivalent:
  1. (1)
    For every nonzero ideal $\mathfrak{q}$ and graded sequence of ideals $\mathfrak{a}_{\bullet}$ on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$, there exists a quasi-monomial valuation $v$ which computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$;
  2. (2)
    $\mathrm{ZVal}_{X}\subseteq\mathrm{QM}_{X}$.
Proof.
$(1)\Rightarrow(2)$: Pick any Zhou valuation $v$ related to some nonzero $\mathfrak{q}$. Then $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1<+\infty$, and there exists a quasi-monomial valuation $w$ computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$ by the assumption. Theorem 4.1 indicates $w=cv$ for some $c\in\mathbf{R}_{+}$, which shows that $v\in\mathrm{QM}_{X}$ is a quasi-monomial valuation.
$(2)\Rightarrow(1)$: For every nonzero ideal $\mathfrak{q}$ and graded sequence of ideals $\mathfrak{a}_{\bullet}$ on $X$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})<+\infty$, by Corollary 4.2, there exists a Zhou valuation $v$ related to $\mathfrak{q}$ that computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. If $\mathrm{ZVal}_{X}\subseteq\mathrm{QM}_{X}$, then $v$ is also quasi-monomial. ∎
Especially, by the solution of the 2-dimensional case of Jonsson–Mustaţă’s conjecture (see Theorem 2.16), Theorem 4.6 induces the following result when $\dim X\leq 2$.
Corollary 4.7.
If $\dim X\leq 2$, then $\mathrm{ZVal}_{X}\subseteq\mathrm{QM}_{X}$.
5. A mixed version of jumping numbers
In this section, we introduce a mixed version of jumping numbers, which is different from the usual so-called mixed jumping numbers (cf. [36]). We also extend the results on jumping numbers from Section 2.3, such as Lemma 2.8, Lemma 2.9, and Lemma 2.12, to their mixed counterparts.
Let $\mathfrak{q},\mathfrak{q}^{\prime}$ be ideals and $\mathfrak{a}$ be a nonzero ideal on $X$. Let $\lambda\in(-\varepsilon_{0},+\infty)$, where $\varepsilon_{0}>0$ so small, and the reason to take the $\varepsilon_{0}$ is that the cases $\lambda<0$ are very important in the later discussions of the present paper. Denote
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})=\mathrm{lct}^{{\mathfrak{q}^{\prime}}^{\lambda}}(\mathfrak{q},\mathfrak{a})\coloneqq\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{a})},$ (5.1)
where the infimum is over all divisors $E$ over $X$ with $\operatorname{ord}_{E}(\mathfrak{a})>0$, and we assume that $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})$ is positive for $\lambda\in(-\varepsilon_{0},+\infty)$. Clearly, this notation coincides with $\mathrm{lct}^{{\mathfrak{q}\cdot\mathfrak{q}^{\prime}}^{k}}(\mathfrak{a})$ when $\lambda=k\in\mathbf{Z}_{\geq 0}$. Moreover, similar to Lemma 2.8, one can get the following lemma.
Lemma 5.1.
If $\pi\colon Y\to X$ is a log resolution of the nonzero ideal $\mathfrak{a}\cdot\mathfrak{q}\cdot\mathfrak{q}^{\prime}$, and if $\mathfrak{a}\cdot\mathcal{O}_{Y}=\mathcal{O}_{Y}(-\sum_{i}\alpha_{i}E_{i})$ and $K_{Y/X}=\sum_{i}\kappa_{i}E_{i}$, then for the ideals $\mathfrak{q}$, $\mathfrak{q}^{\prime}$ and $\lambda\in(-\varepsilon_{0},+\infty)$,
$\begin{split}\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})&=\min_{\alpha_{i}>0}\frac{\kappa_{i}+1+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\alpha_{i}}\\ &=\min_{\operatorname{ord}_{E_{i}}(\mathfrak{a})>0}\frac{A(\operatorname{ord}_{E_{i}})+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E_{i}}(\mathfrak{a})}.\end{split}$
Proof.
Let $E_{i}$ be an irreducible component of $V(\mathfrak{a}\cdot\mathcal{O}_{Y})$. Then we have the following two equations:
  1. (1)
    $\alpha_{i}=\operatorname{ord}_{E_{i}}(\mathfrak{a})$;
  2. (2)
    $A(\operatorname{ord}_{E_{i}})\coloneqq 1+\operatorname{ord}_{E_{i}}(K_{Y/X})=1+\kappa_{i}$.
Therefore, we have the following inequality due to (5.1)
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})\leq\min_{\alpha_{i}>0}\frac{\kappa_{i}+1+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\alpha_{i}}.$
In the following, we prove the inverse inequality. Note that for each divisor $E$, $\operatorname{ord}_{E}(\mathfrak{a})>0$ if and only if $c_{X}(E)\in V(\mathfrak{a})$. For any divisorial valuation $\nu$, we can find a divisor $E$ such that they define the same divisorial valuation $\nu=\nu(E,Y^{\prime})$ up to a scaling, and there is a proper birational morphism $f\colon Y^{\prime}\to Y$ over $X$. So $\operatorname{ord}_{E}(\mathfrak{a})>0$ if and only if $c_{Y}(E)\in V(\mathfrak{a}\cdot\mathcal{O}_{Y})$. Suppose $c_{Y}(E)\in E_{1}\cap\cdots\cap E_{j}$ and $c_{Y}(E)\notin E_{1}\cap\cdots\cap E_{j}\cap E_{j+1}$ with $\alpha_{i}\neq 0$ for all $1\leq i\leq j+1$ for convenience. Suppose further $\mathfrak{q}\cdot\mathcal{O}_{Y}=\mathcal{O}_{Y}(-\sum_{i}t_{i}E_{i})$ and $\mathfrak{q}^{\prime}\cdot\mathcal{O}_{Y}=\mathcal{O}_{Y}(-\sum_{i}s_{i}E_{i})$. Then we have
$\begin{split}A(\operatorname{ord}_{E})&\coloneqq 1+\operatorname{ord}_{E}(K_{Y^{\prime}/X})=1+\operatorname{ord}_{E}(K_{Y^{\prime}/Y})+\operatorname{ord}_{E}(f^{*}(K_{Y/X}))\\ &\geq 1+\operatorname{ord}_{E}(f^{*}E_{1})+\cdots+\operatorname{ord}_{E}(f^{*}E_{j})-1+\operatorname{ord}_{E}\big(f^{*}(\sum_{i}\kappa_{i}E_{i})\big)\\ &=\sum_{1\leq i\leq j}(1+\kappa_{i})\operatorname{ord}_{E}(f^{*}E_{i})\\ &=\sum_{1\leq i\leq j}A(\operatorname{ord}_{E_{i}})\operatorname{ord}_{E}(f^{*}E_{i}).\end{split}$
In addition, we have the following equations:
$\begin{split}\operatorname{ord}_{E}(\mathfrak{a})&=\operatorname{ord}_{E}\Big(f^{*}(\sum_{i}\alpha_{i}E_{i})\Big)=\sum_{1\leq i\leq j}\alpha_{i}\operatorname{ord}_{E}(f^{*}E_{i});\\ \operatorname{ord}_{E}(\mathfrak{q})&=\operatorname{ord}_{E}\Big(f^{*}(\sum_{i}t_{i}E_{i})\Big)=\sum_{1\leq i\leq j}t_{i}\operatorname{ord}_{E}(f^{*}E_{i});\\ \operatorname{ord}_{E}(\mathfrak{q}^{\prime})&=\operatorname{ord}_{E}\Big(f^{*}(\sum_{i}s_{i}E_{i})\Big)=\sum_{1\leq i\leq j}s_{i}\operatorname{ord}_{E}(f^{*}E_{i}).\end{split}$
Combining these relationships, we get
$\begin{split}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{a})}&\geq\min_{1\leq i\leq j}\frac{A(\operatorname{ord}_{E_{i}})+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E_{i}}(\mathfrak{a})}\\ &\geq\min_{\alpha_{i}>0}\frac{\kappa_{i}+1+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\alpha_{i}}.\end{split}$
The proof is done. ∎
Then we can obtain the following corollary which is similar to [31, Lemma 6.7].
Corollary 5.2 (cf. [31, Lemma 6.7]).
For each fixed $\lambda\in(-\varepsilon_{0},+\infty)$, we have
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})=\min_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a})}.$ (5.2)
Suppose that $\mathfrak{a}\neq\mathcal{O}_{X}$ and $(Y,E)$ is a log-smooth pair over $X$ giving a log resolution of $\mathfrak{a}\cdot\mathfrak{q}\cdot\mathfrak{q}^{\prime}$. Then equality in (5.2) is achieved if and only if $v\in\mathrm{QM}(Y,E)$ and $\operatorname{ord}_{E_{i}}$ computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})$ for every irreducible component $E_{i}$ of $E$ for which $v(E_{i})>0$.
Proof.
By Lemma 5.1, it suffices to show that
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})\leq\min_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a})}.$ (5.3)
Let $\pi\colon Y\to X$ be a log resolution of $\mathfrak{a}\cdot\mathfrak{q}\cdot\mathfrak{q}^{\prime}$ with some exceptional effective simple normal crossing divisor $E$. Consider the retraction map $r_{Y,E}\colon\mathrm{Val}_{X}\to\mathrm{QM}(Y,E)$. Then we have $v(*)=r_{Y,E}v(*)$ for $\in\{\mathfrak{a},\mathfrak{q},\mathfrak{q}^{\prime}\}$ by [31, Corollary 4.8], and we have $A(v)\geq A(r_{Y,E}(v))$ by Proposition 2.6. Therefore,
$\min_{v\in\mathrm{QM}(Y,E)}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a})}\leq\min_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a})}.$
Let $\{E_{i}\}$ be the set of smooth irreducible components of $E$. Then for every $v\in\mathrm{QM}(Y,E)$, we have $A(v)=\sum_{i}v(E_{i})\cdot A(\operatorname{ord}_{E_{i}})$ by (2.2), and for all $\in\{\mathfrak{a},\mathfrak{q},\mathfrak{q}^{\prime}\}$ we have $v(*)=\sum_{i}v(E_{i})\cdot\operatorname{ord}_{E_{i}}(*)$ by Proposition 2.3(1). Thus,
$\min_{v\in\mathrm{QM}(Y,E)}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a})}\geq\min_{i}\frac{A(\operatorname{ord}_{E_{i}})+\operatorname{ord}_{E_{i}}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E_{i}}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E_{i}}(\mathfrak{a})}.$
We get (5.3) by applying Lemma 5.1.
Moreover, if $v$ achieves the equality in (5.3), we must have $A(v)=A(r_{Y,E}(v))$ (that is $v\in\mathrm{QM}(Y,E)$ by Proposition 2.6 again) and every $E_{i}$ satisfying $v(E_{i})>0$ computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})$. ∎
Now we can define the mixed version of jumping numbers associated to a graded sequence of ideals in this article as follows. Let $\mathfrak{q},\mathfrak{q}^{\prime}$ be nonzero ideals on $X$. For any nonzero graded sequence $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{m}\}$ of ideals, it can be verified that
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\coloneqq\lim_{m\to\infty}m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})=\sup_{m}m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})\in(0,+\infty]$
is well-defined for any $\lambda\in(-\varepsilon_{0},+\infty)$ with $\varepsilon_{0}$ small enough. Similarly, for any subadditive system of ideals $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{t}\}$ satisfying $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})>0$, we can also define the mixed version of jumping numbers associated to this subadditive sequence by
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})\coloneqq\lim_{t\to\infty}t\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{t})=\inf_{t}t\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{t})\in[0,+\infty)$
for any $\lambda\in(-\varepsilon_{0},+\infty)$. The following results are in the spirit of Lemma 2.9.
Proposition 5.3 ([31, Proposition 2.8]).
If $\mathfrak{b}_{\bullet}$ is a subadditive system of ideals, then
$\begin{split}\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})&=\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}\\ &=\inf_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{b}_{\bullet})},\end{split}$
where $E$ is over all divisors over $X$.
Proof.
Let us first show
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})=\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}.$ (5.4)
On the one hand, we have
$m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{m})\leq\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{m})/m}$
for every divisor $E$ over $X$ by Lemma 5.1. Taking direct limits of $m$, we get
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})\coloneqq\lim_{m\to\infty}m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{m})\leq\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}$
for every $E$, and further gives the inequality
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})\leq\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}.$
On the other hand, for any given $m$, we can find a divisor $E$ over $X$ such that
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{m})=\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{m})}$
by Lemma 5.1 again. Then
$\begin{split}m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{m})&=\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{m})/m}\\ &\geq\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}\\ &\geq\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}\end{split}$
by applying the equations $\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})=\sup_{m}\operatorname{ord}_{E}(\mathfrak{b}_{m})/m$. Thus, we have proven (5.4). Similarly, one can prove
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})=\inf_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{b}_{\bullet})}$
by Corollary 5.2. ∎
Proposition 5.4 ([31, Proposition 2.14]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals and $\mathfrak{q},\mathfrak{q}^{\prime}$ be ideals on $X$. Then $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})$, where $\mathfrak{b}_{\bullet}$ is the subadditive system given by the asymptotic multiplier ideals of $\mathfrak{a}_{\bullet}$.
Proof.
First we have $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})\leq\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{m})$ since $\mathfrak{a}_{m}\subseteq\mathfrak{b}_{m}$. Multiplying by $m$ and letting $m$ go to infinity, it gives $\mathrm{lct}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\leq\mathrm{lct}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})$. Moreover, if $\mathrm{lct}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=+\infty$, then we have $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})=+\infty$ already.
Now, we may assume that $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})<+\infty$. For the reverse inequality, for a given $t>0$, let us choose $m$ divisible enough such that we have $\mathfrak{b}_{t}=\mathcal{J}(\mathfrak{a}_{m}^{t/m})$. Then by [31, Lemma 2.15],
$\frac{\operatorname{ord}_{E}(\mathfrak{a}_{m})}{m}<\frac{\operatorname{ord}_{E}(\mathfrak{b}_{t})}{t}+\frac{A(\operatorname{ord}_{E})}{t}$ (5.5)
for any divisor $E$ over $X$. Pick a divisor $E$ computing $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})>0$. Then we have $A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})>0$ and $\operatorname{ord}_{E}(\mathfrak{a}_{m})>0$. Multiplying both sides of formula (5.5) by $A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})>0$. We get the following inequality
$\begin{split}&\frac{\operatorname{ord}_{E}(\mathfrak{a}_{m})}{m}(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime}))\\ <&\frac{\operatorname{ord}_{E}(\mathfrak{b}_{t})}{t}(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime}))\\ &+\frac{A(\operatorname{ord}_{E})}{t}(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})).\end{split}$
Note that
$\begin{split}&\frac{A(\operatorname{ord}_{E})}{t}\cdot(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime}))\\ =&m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})\cdot\frac{A(\operatorname{ord}_{E})\operatorname{ord}_{E}(\mathfrak{a}_{m})}{mt}\\ \leq&\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\cdot\frac{A(\operatorname{ord}_{E})\operatorname{ord}_{E}(\mathfrak{a}_{m})}{mt}.\end{split}$
Combining the above two inequalities, we get
$\begin{split}&t\cdot{\operatorname{ord}_{E}(\mathfrak{a}_{m})}(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime}))\\ <&m\cdot\operatorname{ord}_{E}(\mathfrak{b}_{t})\big(A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})\big)\\ &+\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\cdot{A(\operatorname{ord}_{E})\cdot\operatorname{ord}_{E}(\mathfrak{a}_{m})}.\end{split}$ (5.6)
We now claim that $\operatorname{ord}_{E}(\mathfrak{b}_{t})\neq 0$ for large enough $t$. Otherwise, we have
$1\leq\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{A(\operatorname{ord}_{E})}<\frac{\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})}{t}$
by (5.6), and this is contradictory to our assumption that $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})<+\infty$. Therefore, dividing by $\operatorname{ord}_{E}(\mathfrak{a}_{m})\cdot\operatorname{ord}_{E}(\mathfrak{b}_{t})$ for (5.6) for large enough $t$, we obtain
$\begin{split}&t\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{t})\\ \leq&\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})/t}\\ \leq&\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\lambda\cdot\operatorname{ord}_{E}(\mathfrak{q}^{\prime})}{\operatorname{ord}_{E}(\mathfrak{a}_{m})/m}+\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\cdot\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})}\\ =&m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})+\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\cdot\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})}\\ \leq&\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})+\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\cdot\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})}.\end{split}$ (5.7)
Note that
$\lim_{t\to\infty}\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})}=\lim_{t\to\infty}\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{t})/t}\cdot\frac{1}{t}=\frac{A(\operatorname{ord}_{E})}{\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})}\lim_{t\to\infty}\frac{1}{t}=0,$
since $\operatorname{ord}_{E}(\mathfrak{b}_{\bullet})\geq\operatorname{ord}_{E}(\mathfrak{b}_{t})/t>0$. Letting $t$ goes to infinity in (5.7), we get
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})\leq\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\qed.$
Corollary 5.5 (cf. [31, Corollary 6.8]).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence of ideals and $\mathfrak{q},\mathfrak{q}^{\prime}$ be ideals on $X$. Then
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\inf_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a}_{\bullet})},$
for each $\lambda\in(-\varepsilon_{0},+\infty)$ such that $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})>0$.
Proof.
By Proposition 5.3 and Proposition 5.4, we have
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})=\inf_{v\in\mathrm{Val}_{X}^{*}}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{b}_{\bullet})}.$
We finish our proof by noting that $v(\mathfrak{b}_{\bullet})=v(\mathfrak{a}_{\bullet})$ whenever $A(v)<\infty$, which is given by [31, Proposition 6.2]. ∎
Lemma 5.6.
Assume that $v\in\mathrm{Val}_{X}^{*}$ computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})<+\infty$. Then
  1. (1)
    $v$ also computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$;
  2. (2)
    $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})=v(\mathfrak{a}_{\bullet})\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$;
  3. (3)
    any $v^{\prime}\in\mathrm{Val}_{X}^{*}$ that computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$ also computes $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$.
Proof.
With the help of Corollary 5.5, this lemma can be proved by the same way of proving Lemma 2.12. ∎
6. Tian functions of valuations
This section is devoted to the proof of Theorem 1.5 (1), following the introduction of the Tian function. Let $\mathfrak{q},\mathfrak{q}^{\prime}$ be ideals on $X$, and $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{m}\}$ a nonzero graded sequence of ideals on $X$. We have the following lemma.
Lemma 6.1.
If $\mathrm{lct}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})<+\infty$, then the function
$(-\varepsilon_{0},+\infty)\ni t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$
is concave and increasing, where $\varepsilon_{0}$ is a sufficiently small positive real number.
Proof.
In fact, for each $t$ and $m\in\mathbf{Z}_{+}$,
$\begin{split}\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})=\inf_{E}\frac{A(\operatorname{ord}_{E})+\operatorname{ord}_{E}(\mathfrak{q})+\operatorname{ord}_{E}(\mathfrak{q}^{\prime})\cdot t}{\operatorname{ord}_{E}(\mathfrak{a}_{m})},\end{split}$
where the infimum is over all divisors $E$ over $X$ with $\operatorname{ord}_{E}(\mathfrak{a}_{m})>0$. Then the function $\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})$ is concave and increasing in $t$ as the infimum of a family of linear functions, for each fixed $m\in\mathbf{Z}_{+}$. It follows that the function $\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\lim\limits_{m\to\infty}m\cdot\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m})$ is also concave and increasing. ∎
The function $t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$ is an algebro-geometric analogue of the so-called Tian function named in [4], so we will also call the function by Tian function. When the graded sequence is associated to a valuation of finite log discrepancy, we have
Proposition 6.2.
Let $v\in\mathrm{Val}_{X}^{*}$ be a valuation with $A(v)<+\infty$. Then for any nonzero ideals $\mathfrak{q}$, $\mathfrak{q}^{\prime}$, the limit
$\lim_{t\to+\infty}\frac{\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})}{t}=v(\mathfrak{q}^{\prime}).$
Proof.
Lemma 6.1 shows the limit
$\mu\coloneqq\lim_{t\to+\infty}\frac{\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})}{t}$
exists. First, we prove $\mu\leq v(\mathfrak{q}^{\prime})$. For each $k\in\mathbf{Z}_{+}$, it holds that
$\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{a}^{v}_{\bullet})=\mathrm{lct}^{\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k}}(\mathfrak{a}^{v}_{\bullet})\leq\frac{A(v)+v(\mathfrak{q})+kv(\mathfrak{q}^{\prime})}{v(\mathfrak{a}^{v}_{\bullet})}=A(v)+v(\mathfrak{q})+kv(\mathfrak{q}^{\prime}),$
which implies $\mu\leq v(\mathfrak{q}^{\prime})$ since $A(v)<+\infty$.
Next, we verify $\mu\geq v(\mathfrak{q}^{\prime})$. We may assume $v(\mathfrak{q}^{\prime})>0$. Set
$\mathfrak{c}_{\bullet}=\{\mathfrak{c}_{m}\}\coloneqq\left\{\mathfrak{q}^{\prime\lceil m/v(\mathfrak{q}^{\prime})\rceil}\right\}_{m=1}^{\infty},$
which is a graded sequence of ideals. Then we have $\mathfrak{c}_{m}\subseteq\mathfrak{a}_{m}^{v}$ for each $m$. Thus, $\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})\leq\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{a}^{v}_{\bullet})<+\infty$ for every nonnegative integer $k$. By the definition of mixed jumping numbers and Corollary 5.5, we get
$\begin{split}\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})&=\inf_{w\in\mathrm{Val}_{X}^{*}}\frac{A(w)+w(\mathfrak{q})+k\cdot w(\mathfrak{q}^{\prime})}{w(\mathfrak{c}_{\bullet})}\\ &=\inf_{w\in\mathrm{Val}_{X}^{*}}\frac{A(w)+w(\mathfrak{q})+k\cdot w(\mathfrak{q}^{\prime})}{w(\mathfrak{q}^{\prime})/v(\mathfrak{q}^{\prime})}\geq k\cdot v(\mathfrak{q}^{\prime}).\end{split}$
It follows that
$\mu=\lim_{k\to\infty}\frac{\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{a}^{v}_{\bullet})}{k}\geq\liminf_{k\to\infty}\frac{\mathrm{lct}(\mathfrak{q},k\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})}{k}\geq v(\mathfrak{q}^{\prime}).$
The proof is complete. ∎
If additionally the valuation computes the jumping number, we can verify that for any nonzero ideal $\mathfrak{q}^{\prime}$ the Tian function is linear on the positive half-axis.
Proposition 6.3.
Let $\mathfrak{q}$ and $\mathfrak{q}^{\prime}$ be two nonzero ideals on $X$, and $v\in\mathrm{Val}_{X}^{*}$ a valuation computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})<+\infty$. Then the Tian function
$t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$
is linear on $0,+\infty)$.
Proof.
Denote $\mathcal{T}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$ for simplicity, where $t\in\mathbf{R}_{\geq 0}$. Due to the concavity of $\mathcal{T}(t)$ given by Lemma 6.1, we only need to prove that for each $k\in\mathbf{Z}_{\geq 0}$, $\mathcal{T}(k)$ is linear in $k$.
Since $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})<+\infty$, we have $A(v)<+\infty$, and for each $w\in\mathrm{Val}_{X}$ with $w\geq v$, it holds that $A(w)+w(\mathfrak{q})\geq A(v)+v(\mathfrak{q})$ by Theorem 2.15. Thus, for such $w$, it also holds that $A(w)+w(\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k})\geq A(v)+v(\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k})$ for each $k$. This shows that $v$ also computes $\mathrm{lct}^{\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k}}(\mathfrak{a}_{\bullet}^{v})$ due to Theorem 2.15, which implies
$\mathcal{T}(k)=\mathrm{lct}^{\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k}}(\mathfrak{a}_{\bullet}^{v})=A(v)+v(\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{k})=A(v)+v(\mathfrak{q})+kv(\mathfrak{q}^{\prime}),\quad\forall k\in\mathbf{Z}_{\geq 0},$
which is linear in $k$. The proof is done. ∎
In fact, the linearity of Tian functions conversely implies that the valuation computes the jumping number. To prove this fact, we first present a technical lemma.
Lemma 6.4.
Let $\mathfrak{q}$ be a nonzero ideal on $X$, and let $v\in\mathrm{Val}_{X}^{*}$ be a valuation with $A(v)<+\infty$. If there exists a nonzero ideal $\mathfrak{q}^{\prime}$ such that $\mathcal{T}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$ is linear in $t\in[t_{0},+\infty)$ for some $t_{0}\geq 0$, then there exists $w\in\mathrm{ZVal}_{X}$ such that the following statements hold:
  1. (1)
    $w\geq v$;
  2. (2)
    $w(\mathfrak{q}^{\prime})=v(\mathfrak{q}^{\prime})$;
  3. (3)
    $\mathrm{lct}(\mathfrak{q},t_{0}\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})=\mathrm{lct}(\mathfrak{q},t_{0}\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{w})=A(w)+w(\mathfrak{q})+t_{0}\cdot w(\mathfrak{q}^{\prime})$.
Proof.
Pick $N\in\mathbf{Z}_{+}$ with $N>t_{0}$. Then by Proposition 3.7 there exists a Zhou valuation $\tilde{w}$ related to $\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{N}$ such that $w\coloneqq\mathcal{T}(N)\tilde{w}\geq v$ and
$\mathrm{lct}(\mathfrak{q},N\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{w})=\mathrm{lct}^{\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{N}}(\mathfrak{a}_{\bullet}^{w})=\mathrm{lct}^{\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{N}}(\mathfrak{a}_{\bullet}^{v})=\mathrm{lct}(\mathfrak{q},N\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})=\mathcal{T}(N).$
Denote $\widetilde{\mathcal{T}}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{w})$, where $t\in[0,+\infty)$. Note $\widetilde{\mathcal{T}}(t)\geq\mathcal{T}(t)$ since $w\geq v$. Since $\widetilde{\mathcal{T}}$ is concave on $(0,+\infty)$ by Lemma 6.1, $\widetilde{\mathcal{T}}(N)=\mathcal{T}(N)$, and $\mathcal{T}$ is linear on $t_{0},+\infty)$, it follows that $\widetilde{\mathcal{T}}\equiv\mathcal{T}$ on $t_{0},+\infty)$. Then since $A(v)<+\infty$, we get $\widetilde{T}(t_{0})=\mathcal{T}(t_{0})$ and
$w(\mathfrak{q}^{\prime})=\lim_{t\to+\infty}\frac{\widetilde{\mathcal{T}}(t)}{t}=\lim_{t\to+\infty}\frac{\mathcal{T}(t)}{t}=v(\mathfrak{q}^{\prime}),$
according to Proposition 6.2.
What remains to be proved for $w$ is $\mathrm{lct}(\mathfrak{q},t_{0}\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{w})=A(w)+w(\mathfrak{q})+t_{0}\cdot w(\mathfrak{q}^{\prime})$. Since $\tilde{w}$ is a Zhou valuation related to $\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{N}$, due to Theorem 4.1 we have
$\widetilde{\mathcal{T}}(N)=\mathcal{T}(N)=A(w)+w(\mathfrak{q}\cdot{\mathfrak{q}^{\prime}}^{N})=A(w)+w(\mathfrak{q})+N\cdot w(\mathfrak{q}^{\prime}).$
Thus,
$\begin{split}\widetilde{\mathcal{T}}(t_{0})=\mathcal{T}(t_{0})=&\mathcal{T}(N)-(N-t_{0})\cdot v(\mathfrak{q}^{\prime})\\ =&A(w)+w(\mathfrak{q})+N\cdot w(\mathfrak{q}^{\prime})-(N-t_{0})\cdot w(\mathfrak{q}^{\prime})\\ =&A(w)+w(\mathfrak{q})+t_{0}\cdot w(\mathfrak{q}^{\prime}).\end{split}$
The proof is complete. ∎
Now we prove the following proposition, which is a converse of Proposition 6.3. And this completes the proof of Theorem 1.5 (1).
Proposition 6.5.
Let $\mathfrak{q}$ be a nonzero ideal on $X$ and $v\in\mathrm{Val}_{X}^{*}$ with $A(v)<+\infty$. If the function
$t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$
is linear in $t\in[0,+\infty)$ for every nonzero ideal $\mathfrak{q}^{\prime}$ on $X$, then $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})$.
Proof.
Assume that $v$ does not compute $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})\in(0,+\infty)$, i.e.,
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})<A(v)+v(\mathfrak{q}),$
and we prove the desired result by contradiction.
Without loss of generality, we assume $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$. By the definition of log-discrepancy, there exists a log-smooth pair $(Y,E)$ over $X$ such that $v^{\prime}\coloneqq r_{Y,E}(v)$ satisfies
$A(v^{\prime})+v^{\prime}(\mathfrak{q})=A(v)+v(\mathfrak{q})-\varepsilon>1.$
After replacing $X$ by an open neighborhood of the center $c_{X}(v^{\prime})$ of $v^{\prime}$ on $X$, we may assume that $X=\operatorname{Spec}R$ is affine. Since $v^{\prime}=r_{Y,E}(v)$, there exist algebraic local coordinates $y_{1},\ldots,y_{N}$ at $c_{Y}(v^{\prime})$ with respect to which $v^{\prime}$ is monomial and $E_{i}=\{y_{i}=0\}$, $1\leq i\leq N$. Write $y_{i}=a_{i}/b_{i}$ with $a_{i},b_{i}\in R$ nonzero. Set
$\mathfrak{q}^{\prime}=\mathfrak{q}\cdot(b_{1}\cdots b_{N}).$
Then according to Lemma 6.4, we can find some $w\in\mathrm{ZVal}_{X}$ such that $w\geq v$, $w(\mathfrak{q}^{\prime})=v(\mathfrak{q}^{\prime})$, and
$1=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=A(w)+w(\mathfrak{q}).$
It follows that $w(\mathfrak{q})=v(\mathfrak{q})\geq v^{\prime}(\mathfrak{q})$, and $w(b_{i})=v(b_{i})$ for $1\leq i\leq N$. Thus,
$w(E_{i})=w(a_{i})-v(b_{i})\geq v(a_{i})-v(b_{i})=v(E_{i})=v^{\prime}(E_{i}),\quad 1\leq i\leq N.$
Then $A(w)\geq A(r_{Y,E}(w))\geq A(v^{\prime})$ due to (2.2). However,
$1=A(w)+w(\mathfrak{q})\geq A(v^{\prime})+v^{\prime}(\mathfrak{q})>1,$
which is a contradiction. The proof of that $v$ computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})$ is done. ∎
7. Enlarging a graded sequence
This section provides a natural generalization of [31, Section 7.4] to the context of mixed jumping numbers. In addition, it serves as preparation for the proof of Theorem 1.5 (2).
Let $\mathfrak{a}_{\bullet}$ be a graded sequence, and $\mathfrak{q}^{\prime}$ be a nonzero ideal on $X$. Note that for every $\beta>0$, if we set
$\mathfrak{c}_{m}\coloneqq\sum_{i=0}^{m}\mathfrak{a}_{i}\cdot{\mathfrak{q}^{\prime}}^{\lceil\beta(m-i)\rceil},\quad\forall\,m\in\mathbf{Z}_{\geq 0},$
then $\mathfrak{c}_{\bullet}=\{\mathfrak{c}_{m}\}$ is a graded sequence such that $\mathfrak{c}_{m}\supseteq\mathfrak{a}_{m}\cup{\mathfrak{q}^{\prime}}^{\lceil\beta m\rceil}$ for each $m\in\mathbf{Z}_{\geq 0}$. This is a way to enlarge a graded sequence (see [31, Section 7.4]). We first establish a result for mixed jumping numbers, which is the key lemma to prove the differentiability part of Theorem 1.5 (2).
Lemma 7.1.
Let $\mathfrak{q},\mathfrak{q}^{\prime}$ be two nonzero ideals, and $\mathfrak{a}_{\bullet}$ a graded sequence of ideals. Let $\beta>0$ and denote by $\mathfrak{c}_{\bullet}=\{\mathfrak{c}_{m}\}$ the graded sequence given by
$\mathfrak{c}_{m}=\sum_{i=0}^{m}\mathfrak{a}_{i}\cdot{\mathfrak{q}^{\prime}}^{\lceil\beta(m-i)\rceil},\quad m\in\mathbf{Z}_{\geq 0}.$
Given $\lambda\in\mathbf{R}$, if
$0<\mathrm{lct}(\mathfrak{q},(\lambda-\beta d)\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})+d\leq\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})<+\infty$
for every sufficiently small $d>0$, then $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})=\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$.
Proof.
Since $\mathfrak{c}_{m}\supseteq\mathfrak{a}_{m}$ for each $m$, we have $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})\geq\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$. For each $v\in\mathrm{Val}_{X}^{*}$,
$\begin{split}v(\mathfrak{c}_{m})&=v\left(\sum_{i=0}^{m}\mathfrak{a}_{i}\cdot{\mathfrak{q}^{\prime}}^{\lceil\beta(m-i)\rceil}\right)\\ &=\min_{i}\big\{v(\mathfrak{a}_{i})+\left\lceil\beta(m-i)\right\rceil\cdot v(\mathfrak{q}^{\prime})\big\}\\ &=\min\big\{\beta\cdot v(\mathfrak{q}^{\prime})m,v(\mathfrak{a}_{m})\big\}+O(1),\end{split}$
which shows
$v(\mathfrak{c}_{\bullet})=\min\big\{\beta\cdot v(\mathfrak{q}^{\prime}),v(\mathfrak{a}_{\bullet})\big\}.$ (7.1)
For every sufficiently small $d>0$, pick any $w\in\mathrm{Val}_{X}^{*}$ such that $w(\mathfrak{a}_{\bullet})=1$ and
$\mathrm{lct}(\mathfrak{q},(\lambda-\beta d)\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=A(w)+w(\mathfrak{q})+(\lambda-\beta d)\cdot w(\mathfrak{q}^{\prime})-\delta(d)\cdot d,$
where $\delta(d)\to 0^{+}$ when $d\to 0^{+}$. Then
$\begin{split}&A(w)+w(\mathfrak{q})+(\lambda-\beta d)\cdot w(\mathfrak{q}^{\prime})-\delta(d)\cdot d\\ =&\mathrm{lct}(\mathfrak{q},(\lambda-\beta d)\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\\ \leq&\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})-d\\ \leq&\frac{A(w)+w(\mathfrak{q})+\lambda\cdot w(\mathfrak{q}^{\prime})}{w(\mathfrak{a}_{\bullet})}-d\\ =&A(w)+w(\mathfrak{q})+\lambda\cdot w(\mathfrak{q}^{\prime})-d,\end{split}$
yielding $\beta\cdot w(\mathfrak{q}^{\prime})\geq 1-\delta(d)$. Combining with (7.1), we get $w(\mathfrak{c}_{\bullet})\geq 1-\delta(d)$. It follows that
$\begin{split}\mathrm{lct}(\mathfrak{q},(\lambda-\beta d)\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})&\leq\frac{A(w)+w(\mathfrak{q})+(\lambda-\beta d)\cdot w(\mathfrak{q}^{\prime})}{w(\mathfrak{c}_{\bullet})}\\ &\leq\frac{A(w)+w(\mathfrak{q})+(\lambda-\beta d)\cdot w(\mathfrak{q}^{\prime})}{1-\delta(d)}\\ &=\frac{\mathrm{lct}(\mathfrak{q},(\lambda-\beta d)\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})+\delta(d)\cdot d}{1-\delta(d)}.\end{split}$
Let $d\to 0^{+}$, and by the continuity of Tian functions (see Lemma 6.1), we obtain $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})\leq\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$.
Consequently, $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{c}_{\bullet})=\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})$. ∎
Moreover, for Zhou valuations, we have the following
Lemma 7.2.
Let $v$ be a Zhou valuation related to a nonzero ideal $\mathfrak{q}$, and $\mathfrak{q}^{\prime}$ a nonzero ideal. Assume that a graded sequence $\mathfrak{c}_{\bullet}=\{\mathfrak{c}_{m}\}$ of ideals satisfying
$\mathfrak{c}_{m}=\sum_{i=0}^{m}\mathfrak{a}^{v}_{i}\cdot{\mathfrak{q}^{\prime}}^{\lceil\beta(m-i)\rceil},\quad m\in\mathbf{Z}_{\geq 0},$
where $\beta$ is a positive constant. Then we have $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})=1$ if $\beta\geq 1/v(\mathfrak{q}^{\prime})$, and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})>1$ if $\beta<1/v(\mathfrak{q}^{\prime})$.
Proof.
Note that we have $\mathfrak{c}_{m}\supseteq\mathfrak{a}_{m}^{v}\cup{\mathfrak{q}^{\prime}}^{\lceil\beta m\rceil}$ for all $m\in\mathbf{Z}_{\geq 0}$. Since $\mathfrak{a}^{v}_{m}\subseteq\mathfrak{c}_{m}$ for each $m$, we have $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})\geq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{v}_{\bullet})=1$.
If $\beta\geq 1/v(\mathfrak{q}^{\prime})$, then $v\left({\mathfrak{q}^{\prime}}^{\lceil\beta m\rceil}\right)=\lceil\beta m\rceil\cdot v(\mathfrak{q}^{\prime})\geq m$, which implies ${\mathfrak{q}^{\prime}}^{\lceil\beta m\rceil}\subseteq\mathfrak{a}_{m}^{v}$ and thus $\mathfrak{a}_{m}^{v}=\mathfrak{c}_{m}$ for each $m$. It is clear that $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})=1$ for this situation.
Next we assume $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})=1$. Then Theorem 3.3 shows that there exists a Zhou valuation $w$ related to $\mathfrak{q}$ satisfying $\mathfrak{c}_{m}\subseteq\mathfrak{a}_{m}^{w}$ for every $m$. This implies $\mathfrak{a}_{m}^{v}\subseteq\mathfrak{a}_{m}^{w}$ for each $m$, and then $w\geq v$. The fact that $v$ is a Zhou valuation related to $\mathfrak{q}$ indicates $w=v$. We get $\mathfrak{a}_{m}^{v}=\mathfrak{c}_{m}$ for each $m$. It follows that $v\left({\mathfrak{q}^{\prime}}^{\lceil\beta m\rceil}\right)\geq m$, which gives $\beta\geq 1/v(\mathfrak{q}^{\prime})$. Consequently, we have $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})>1$ if $\beta<1/v(\mathfrak{q}^{\prime})$. ∎
8. Tian functions of Zhou valuations
In this section, we prove Theorem 1.5 (2) by enlarging graded sequences, as in the previous section. First, we prove that a valuation being a Zhou valuation will imply the linearity on $0,+\infty)$ and the differentiability at $t=0$ of Tian functions.
Theorem 8.1.
Let $\mathfrak{q}$ and $\mathfrak{q}^{\prime}$ be two nonzero ideals on $X$, and $v\in\mathrm{Val}_{X}^{*}$ a Zhou valuation related to $\mathfrak{q}$. Then the function
$(-\varepsilon_{0},+\infty)\ni t\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$
is linear on $0,+\infty)$, and differentiable at $t=0$.
Proof of Theorem 8.1.
Denote $\mathcal{T}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$ for $t\in(-\varepsilon_{0},+\infty)$. By Theorem 4.1 and Proposition 6.3, the linearity of $\mathcal{T}(t)$ on $0,+\infty)$ holds true. In the following we prove that $\mathcal{T}(t)$ is differentiable at $t=0$.
Due to the monotonicity and concavity of $\mathcal{T}(t)$, we have that $\mathcal{T}^{\prime}_{+}(0)$ and $\mathcal{T}^{\prime}_{-}(0)$ exist, and $0\leq\mathcal{T}^{\prime}_{+}(0)\leq\mathcal{T}^{\prime}_{-}(0)$. It suffices to prove $\mathcal{T}^{\prime}_{+}(0)\geq\mathcal{T}^{\prime}_{-}(0)$. We may assume $\mathcal{T}^{\prime}_{-}(0)>0$. Denote $\gamma\coloneqq 1/\mathcal{T}^{\prime}_{-}(0)$. Then for every sufficiently small $d>0$, we have
$\mathcal{T}\left(-\gamma d\right)+d\leq\mathcal{T}(0).$
Set $\mathfrak{c}_{\bullet}=\{\mathfrak{c}_{m}\}$, where
$\mathfrak{c}_{m}=\sum_{i=0}^{m}\mathfrak{a}^{v}_{i}\cdot{\mathfrak{q}^{\prime}}^{\lceil\gamma(m-i)\rceil},\quad m\in\mathbf{Z}_{\geq 0}.$
Now Proposition 3.5 and Lemma 7.1 yield that $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{c}_{\bullet})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$, and it follows from Lemma 7.2 that
$\gamma\geq 1/v(\mathfrak{q}^{\prime})\Rightarrow\mathcal{T}^{\prime}_{-}(0)\leq v(\mathfrak{q}^{\prime}).$
Proposition 6.3 shows $\mathcal{T}^{\prime}_{+}(0)=v(\mathfrak{q}^{\prime})$. Thus, we get $\mathcal{T}^{\prime}_{+}(0)\geq\mathcal{T}^{\prime}_{-}(0)$, which completes the proof of that $\mathcal{T}(t)$ is differentiable at $t=0$. ∎
Conversely, we have the following result.
Proposition 8.2.
Let $\mathfrak{q}$ be a nonzero ideal on $X$. If $v\in\mathrm{Val}_{X}^{*}$ with $A(v)<+\infty$ satisfying that the function
$\begin{split}\mathcal{T}_{\mathfrak{q}^{\prime}}(t)\colon(-\varepsilon_{\mathfrak{q}^{\prime}},+\infty)&\longrightarrow\mathbf{R}_{\geq 0}\\ t&\longmapsto\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})\end{split}$
is differentiable at $t=0$ for every nonzero ideal $\mathfrak{q}^{\prime}$ on $X$, then
$\mathfrak{q}^{\prime}\longmapsto\frac{\mathrm{d}\mathcal{T}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}$
is a valuation on $X$ belonging to $\mathrm{ZVal}_{X}$. If additionally assume that $\mathcal{T}_{\mathfrak{q}^{\prime}}(t)$ is linear on $0,+\infty)$ for every ideal $\mathfrak{q}^{\prime}$ on $X$ and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$, then $v$ is a Zhou valuation related to $\mathfrak{q}$.
Proof.
Let $w\in\mathrm{ZVal}_{X}$ with $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})$ and $w\geq v$, given by Proposition 3.7, and the associated Zhou valuation of $w$ is related to $\mathfrak{q}$. Denote
$\widetilde{\mathcal{T}}_{\mathfrak{q}^{\prime}}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{w}),$
for each nonzero ideal $\mathfrak{q}^{\prime}$ on $X$. Then $\widetilde{\mathcal{T}}_{\mathfrak{q}^{\prime}}\geq\mathcal{T}_{\mathfrak{q}^{\prime}}$, and by (a variation of) Theorem 8.1, $\widetilde{\mathcal{T}}_{\mathfrak{q}^{\prime}}$ is differentiable at $t=0$ with
$\frac{\mathrm{d}\widetilde{\mathcal{T}}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}=w(\mathfrak{q}^{\prime}).$
Since $\mathcal{T}_{\mathfrak{q}^{\prime}}$ is concave and differentiable at $t=0$, it must hold that
$\frac{\mathrm{d}\mathcal{T}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}=\frac{\mathrm{d}\widetilde{\mathcal{T}}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}=w(\mathfrak{q}^{\prime}).$
It follows that $\mathfrak{q}^{\prime}\mapsto\frac{\mathrm{d}\mathcal{T}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}$ is a Zhou valuation (equals to $w$). If additionally we have that $\mathcal{T}_{\mathfrak{q}^{\prime}}(t)$ is linear on $0,+\infty)$ for every ideal $\mathfrak{q}^{\prime}$ on $X$ and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$, then by Proposition 6.2,
$\frac{\mathrm{d}\mathcal{T}_{\mathfrak{q}^{\prime}}}{\mathrm{d}t}\Big|_{t=0}=v(\mathfrak{q}^{\prime})\quad\text{and}\quad\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1,$
and thus $w$ can be chosen as a Zhou valuation related to $\mathfrak{q}$. Consequently, $v=w$ is also a Zhou valuation related to $\mathfrak{q}$. ∎
9. Denseness of the cone of Zhou valuations
We demonstrate that the cone of Zhou valuations is dense in the cone of valuations. A similar statement can be found in [5, Section 1.4], and our proof here is almost the same of there.
Lemma 9.1.
Let $v\in\mathrm{Val}_{X}^{*}$ with $A(v)<+\infty$. Then for every nonzero ideal $\mathfrak{q}$ on $X$,
$v(\mathfrak{q})=\inf\big\{\tilde{v}(\mathfrak{q})\colon\tilde{v}\in\mathrm{ZVal}_{X},\ \tilde{v}\geq v\big\}.$
Proof.
It suffices to verify that
$v(\mathfrak{q})\geq\inf\big\{\tilde{v}(\mathfrak{q})\colon\tilde{v}\in\mathrm{ZVal}_{X},\ \tilde{v}\geq v\big\}.$
For every $k\in\mathbf{Z}_{+}$, we have $\gamma(k)\coloneqq\mathrm{lct}^{\mathfrak{q}^{k}}(\mathfrak{a}_{\bullet}^{v})<+\infty$, since $A(v)<+\infty$. Then for every given $k$ we have $\mathrm{lct}^{\mathfrak{q}^{k}}(\mathfrak{a}_{\bullet}^{v/\gamma(k)})=1$, and there exists a Zhou valuation $\tilde{v}_{k}$ related to $\mathfrak{q}^{k}$ such that
  1. (1)
    $\frac{v}{\gamma(k)}\leq\tilde{v}_{k}$ by Lemma 3.1 and Corollary 3.4;
  2. (2)
    $\tilde{v}_{k}(\mathfrak{q}^{k})=1-A(\tilde{v}_{k})\leq 1$ by Theorem 4.1.
Let $\hat{v}_{k}\coloneqq\gamma(k)\cdot\tilde{v}_{k}$. Then $\hat{v}_{k}\in\mathrm{ZVal}_{X}$ with $\hat{v}_{k}\geq v$, and
$v(\mathfrak{q})\leq\hat{v}_{k}(\mathfrak{q})=\frac{\gamma(k)}{k}\cdot\tilde{v}_{k}(\mathfrak{q}^{k})\leq\frac{\mathrm{lct}^{\mathfrak{q}^{k}}(\mathfrak{a}_{\bullet}^{v})}{k}\leq v(\mathfrak{q})+\frac{A(v)}{k},$
where the last inequality is given by Lemma 2.10. Due to $A(v)<+\infty$, we obtain
[…]
and finish the proof. ∎
Theorem 9.2.
The cone of the Zhou valuations $\mathrm{ZVal}_{X}$ on $X$ is dense in $\mathrm{Val}_{X}$.
Proof.
Note that for every quasi-monomial valuation $v$ we have $A(v)<+\infty$. Therefore, by Proposition 2.4(1), it suffices to show that for every $v\in\mathrm{Val}_{X}^{*}$ with $A(v)<+\infty$, for any finite fixed non-trivial and nonzero ideals $\mathfrak{q}_{1},\cdots,\mathfrak{q}_{r}$ ($r\in\mathbf{Z}_{+}$), and for every given integer $n\in\mathbf{Z}_{+}$, there exists a valuation $\tilde{v}\in\mathrm{ZVal}_{X}$ such that
$\max_{1\leq i\leq r}|\tilde{v}(\mathfrak{q}_{i})-v(\mathfrak{q}_{i})|<\frac{1}{n}.$
Set $\mathfrak{q}\coloneqq\mathfrak{q}_{1}\cdots\mathfrak{q}_{r}$. Then according to Lemma 9.1, for any $\epsilon>0$, we can find a valuation $\tilde{v}\in\mathrm{ZVal}_{X}$ with ${v}\leq\tilde{v}$ and $\tilde{v}(\mathfrak{q})<v(\mathfrak{q})+\epsilon$. Then we can choose $\epsilon$ to be small enough so that $|\tilde{v}(\mathfrak{q}_{i})-v(\mathfrak{q}_{i})|<\frac{1}{n}$ for all $i\in\{1,\cdots,r\}$. ∎
10. Singularities of graded sequences and Zhou valuations
One important application of Zhou weights in [4] is to characterize the singularities of plurisubharmonic functions, which recovers and generalizes a part of the main result of [10], while the Zhou valuations are also used to characterize the division relation of germs of holomorphic functions in [4]. In this section, we show that the algebraic Zhou valuations can be used to characterize the singularities of graded sequences of ideals. Our result is as follows:
Theorem 10.1.
Let $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{m}\}_{m\geq 1}$, $\mathfrak{a}^{\prime}_{\bullet}=\{\mathfrak{a}^{\prime}_{m}\}_{m\geq 1}$ be two nonzero graded sequences of ideals on $X$. Then the following statements are equivalent:
  1. (1)
    $\mathcal{J}(t\cdot\mathfrak{a}_{\bullet})\subseteq\mathcal{J}(t\cdot\mathfrak{a}^{\prime}_{\bullet})$ for every $t>0$.
  2. (2)
    $v(\mathfrak{a}_{\bullet})\geq v(\mathfrak{a}^{\prime}_{\bullet})$ for every $v\in\mathrm{Val}_{X}$ with $A(v)<+\infty$.
  3. (3)
    $v(\mathfrak{a}_{\bullet})\geq v(\mathfrak{a}^{\prime}_{\bullet})$ for every $v\in\mathrm{ZVal}_{X}$.
Proof.
Obviously, we have (2) $\Rightarrow$ (3). Then it suffices to prove (1) $\Rightarrow$ (2) and (3) $\Rightarrow$ (1).
(1) $\Rightarrow$ (2): Without loss of generality, we assume $v(\mathfrak{a}^{\prime}_{\bullet})>0$. Fix $k\in\mathbf{Z}_{\geq 0}$ first. Denote $\mathfrak{q}_{k}\coloneqq\mathcal{J}(k\cdot\mathfrak{a}_{\bullet})$. Since $\mathfrak{q}_{k}\subseteq\mathcal{J}(k\cdot\mathfrak{a}^{\prime}_{\bullet})$, we have
$\frac{A(v)+v(\mathfrak{q}_{k})}{v(\mathfrak{a}^{\prime}_{\bullet})}\geq\mathrm{lct}^{\mathfrak{q}_{k}}(\mathfrak{a}^{\prime}_{\bullet})>k.$ (10.1)
On the other hand, we can find an integer $p_{k}>0$ such that
$\mathcal{J}(k\cdot\mathfrak{a}_{\bullet})=\mathcal{J}\Big(\frac{k}{pk}\cdot\mathfrak{a}_{pk}\Big)=\mathcal{J}\Big(\frac{1}{p}\cdot\mathfrak{a}_{pk}\Big)$
for all $p\geq p_{k}$ (cf. [36, Prop. 11.1.18]). The subadditivity property of multiplier ideals (see [31, Theorem A.2] or [36, Theorem 9.5.17]) implies
$\mathfrak{q}_{k}^{p}=\mathcal{J}\Big(\frac{1}{p}\cdot\mathfrak{a}_{pk}\Big)^{p}\supseteq\mathcal{J}\Big(\frac{p}{p}\cdot\mathfrak{a}_{pk}\Big)=\mathcal{J}(\mathfrak{a}_{pk})\supseteq\mathfrak{a}_{pk},$
yielding that
$v(\mathfrak{q}_{k})\leq\frac{1}{p}v(\mathfrak{a}_{pk})\quad\text{for}\quad\forall p\geq p_{k}.$
Combining with (10.1) we get
$v(\mathfrak{a}^{\prime}_{\bullet})<\frac{A(v)+v(\mathfrak{q}_{k})}{k}\leq\frac{A(v)}{k}+\frac{1}{pk}v(\mathfrak{a}_{pk}).$
Let $p\to+\infty$, and then it follows that
$v(\mathfrak{a}^{\prime}_{\bullet})\leq\frac{A(v)}{k}+v(\mathfrak{a}_{\bullet}).$
Finally, we let $k\to+\infty$, and thus getting $v(\mathfrak{a}^{\prime}_{\bullet})\leq v(\mathfrak{a}_{\bullet})$.
(3) $\Rightarrow$ (1): If $(1)$ is not true, then there exist $t_{0}>0$ and a nonzero ideal $\mathfrak{q}$ on $X$ such that $\mathfrak{q}\subseteq\mathcal{J}(t_{0}\cdot\mathfrak{a}_{\bullet})$ and $\mathfrak{q}\not\subseteq\mathcal{J}(t_{0}\cdot\mathfrak{a}^{\prime}_{\bullet})$. Thus,
$t_{1}\coloneqq\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}^{\prime}_{\bullet})\leq t_{0}<\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}).$
Since $\mathfrak{a}_{\bullet}^{\prime}$ is nonzero, we have $t_{1}>0$. By Corollary 4.2, there exists a Zhou valuation related to $\mathfrak{q}$ with $v(\mathfrak{a}_{\bullet}^{\prime})=1/t_{1}$, computing $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{\prime})$. Then it follows from the assumption of (3) that
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})\leq\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}\leq\frac{1}{v(\mathfrak{a}_{\bullet}^{\prime})}=t_{1},$
which contradicts to $t_{1}<\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})$. Thus, (3) $\Rightarrow$ (1).
The proof is complete. ∎
The following proposition is a characterization of the asymptotic multiplier ideals via Zhou valuations which is an analogue of [4, Corollary 1.17].
Proposition 10.2.
For every nonzero ideal $\mathfrak{q}\subseteq\mathcal{O}_{X}$, $\mathfrak{a}_{\bullet}$ a graded sequence of ideals on $X$, and $\lambda>0$, we have $\mathfrak{q}\subseteq\mathcal{J}(\lambda\cdot\mathfrak{a}_{\bullet})$ if and only if $v(\mathfrak{a}_{\bullet})<\frac{1}{\lambda}$ for every Zhou valuation $v$ related to $\mathfrak{q}$ on $X$.
Proof.
This proposition holds since
$\begin{split}\mathfrak{q}\subseteq\mathcal{J}(\lambda\cdot\mathfrak{a}_{\bullet})\quad\Longleftrightarrow&\quad\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet})>\lambda\\ \Longleftrightarrow&\quad\min_{v\in\mathrm{ZVal}_{X}(\mathfrak{q})}\frac{A(v)+v(\mathfrak{q})}{v(\mathfrak{a}_{\bullet})}>\lambda\\ \Longleftrightarrow&\quad\min_{v\in\mathrm{ZVal}_{X}(\mathfrak{q})}\frac{1}{v(\mathfrak{a}_{\bullet})}>\lambda\\ \Longleftrightarrow&\quad\max_{v\in\mathrm{ZVal}_{X}(\mathfrak{q})}v(\mathfrak{a}_{\bullet})<\frac{1}{\lambda},\end{split}$
where we recall that $\mathrm{ZVal}_{X}(\mathfrak{q})$ denotes the set of Zhou valuations related to $\mathfrak{q}$ on $X$. ∎
11. Analytic correspondence
Since the ideal of algebraic Zhou valuations actually comes from the analytic Zhou valuations, it is natural to consider whether one can reduce the algebraic Zhou valuations to the original analytic Zhou valuations on some special schemes.
In [32], the authors gave the ways to link the objects in algebraic and analytic settings. We recall some constructions and observations from [32]. Let $U$ be a complex manifold, and let $V$ be the germ of a complex submanifold at a point $x$ in $U$. Denote by $\mathcal{O}_{x,V}$ the localization of $\mathcal{O}_{x}$ along the ideal $I_{V}$, where $\mathcal{O}_{x}$ denotes the ring of germs of holomorphic functions at $x$. Then $\mathcal{O}_{x,V}$ is an excellent regular domain with maximal ideal $\mathfrak{m}_{x,V}$. Set
$\mathrm{lct}_{x,V}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})\coloneqq\mathrm{lct}(\mathfrak{q}\cdot\mathcal{O}_{x,V},\lambda\cdot(\mathfrak{q}^{\prime}\cdot\mathcal{O}_{x,V});\mathfrak{a}_{\bullet}\cdot\mathcal{O}_{x,V}),$
and
$\mathrm{lct}_{x,V}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})\coloneqq\mathrm{lct}(\mathfrak{q}\cdot\mathcal{O}_{x,V},\lambda\cdot(\mathfrak{q}^{\prime}\cdot\mathcal{O}_{x,V});\mathfrak{b}_{\bullet}\cdot\mathcal{O}_{x,V}),$
where $\mathfrak{q},\mathfrak{q}^{\prime}$ are ideals of $\mathcal{O}_{x}$, $\mathfrak{a}_{\bullet}$ is a graded sequence of ideals on $X$, and $\mathfrak{b}_{\bullet}$ is a subadditive system of ideals on $X$. Then
$\mathrm{lct}_{x,V}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet})=\inf_{v}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{a}_{\bullet})}$
and
$\mathrm{lct}_{x,V}(\mathfrak{q},\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet})=\inf_{v}\frac{A(v)+v(\mathfrak{q})+\lambda\cdot v(\mathfrak{q}^{\prime})}{v(\mathfrak{b}_{\bullet})},$
where the infimums are over all (quasi-monomial, or Zhou) valuations of $\mathcal{O}_{x,V}$ by Theorem 9.2. In this section, we only care about the case $V=\{x\}$.
Let $\varphi$ be a plurisubharmonic (psh for short) function on a complex manifold $U$. Then the multiplier ideal sheaf $\mathcal{J}(\varphi)$ is a sheaf on $U$ such that at any point $x\in U$ the stalk is defined by
$\mathcal{J}(\varphi)_{x}\coloneqq\big\{f\in\mathcal{O}_{x}\colon|f|^{2}e^{-2\varphi}\text{ is locally integrable at }x\big\}.$ (11.1)
The sheaf $\mathcal{J}(\varphi)$ is a coherent sheaf ([38, 39]). For non-zero ideals $\mathfrak{q},\mathfrak{q}^{\prime}$ on $U$ and $\lambda\in\mathbf{R}$, define the mixed jumping number of $\varphi$ at $x$ relative to $\mathfrak{q},\mathfrak{q}^{\prime}$ and $\lambda$ by
$c^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}_{x}(\varphi)\coloneqq\sup\big\{c\geq 0\colon|\mathfrak{q}|^{2}|\mathfrak{q}^{\prime}|^{2\lambda}e^{-2c\varphi}\text{ is locally integrable at }x\big\},$
where it is the classical jumping number if $\mathfrak{q}^{\prime}=\mathcal{O}_{U}$ or $\lambda=0$.
The following proposition is [32, Proposition 3.12] after some slight changes.
Proposition 11.1 ([32]).
Let $x$ be any point in $U$, and $V$ be the germ of a proper complex submanifold at $x$. Let $\varphi$ be a psh function on $U$, and set
$\mathfrak{b}_{t}\coloneqq\mathcal{J}(t\varphi),\ t>0,$
which is a subadditive sequence of ideals on $U$. The following statements hold.
  1. (1)
    For nonzero ideals $\mathfrak{q},\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{x}$ and $\lambda\in\mathbf{R}$, we have
    $c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi)=\mathrm{lct}_{x}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet}).$
  2. (2)
    The subadditive sequence $\mathfrak{b}_{\bullet}\cdot\mathcal{O}_{x,V}$ has controlled growth.
  3. (3)
    If $\varphi=\log|\mathfrak{a}|+O(1)$ near $o$ for some nonzero ideal $\mathfrak{a}\subseteq\mathcal{O}_{x}$, then $\mathfrak{b}_{t}=\mathcal{J}(t\cdot\mathfrak{a})$ for every $t>0$. Therefore, we have $c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi)=\mathrm{lct}_{x}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a})$ for this case.
Proof.
Here we follow the arguments from [32, Proposition 3.12], where the statement (2) has been proved.
To prove (1), we use Demailly’s approximation theorem (see [12, 14]). Take a pseudoconvex domain $B\subseteq U$ containing $x$ and for every $p\geq 1$, define
$\varphi_{p}\coloneqq\frac{1}{p}\sup\left\{\log|f|\colon\int_{B}|f|^{2}e^{-2p\varphi}\leq 1\right\}.$
Then $\varphi_{p}$ is psh on $B$. By the Ohsawa-Takegoshi $L^{2}$ extension theorem ([40]), there exists a constant $C$ not depending on the choices of $\varphi$ and $p$ such that
$\varphi\leq\varphi_{p}+\frac{C}{p}$
on $B$. Therefore, for any nonzero ideals $\mathfrak{q}$ and $\mathfrak{q}^{\prime}$ of $\mathcal{O}_{x}$, we have
$\begin{split}(p\cdot\mathrm{lct}_{x}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{p}))^{-1}&=c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi_{p})^{-1}\leq c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi)^{-1}\\ &\leq c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi_{p})^{-1}+\frac{1}{p}\\ &=(p\cdot\mathrm{lct}_{x}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{p}))^{-1}+\frac{1}{p}\end{split}$
for every $p\geq 1$, where the two equalities follow from [32, (3.12)] in the general case. Letting $p$ go to infinity and applying the equation before Proposition 5.3, we get
$c_{x}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\varphi)=\mathrm{lct}_{x}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{b}_{\bullet}).$
It remains to prove (3). This is given by the following
$\mathfrak{b}_{t}\coloneqq\mathcal{J}(t\varphi)=\mathcal{J}(t\log|\mathfrak{a}|)=\mathcal{J}(\log|\mathfrak{a}^{t}|)=\mathcal{J}(t\cdot\mathfrak{a}),$
where the last equation follows from [16, Proposition 1.7]. Finally, we finish our proof by applying Proposition 11.1(1) and Proposition 5.4. ∎
11.1. The analytic correspondence of algebraic Zhou valuations
Let $o$ be the origin in $\mathbf{C}^{n}$. Denote by $\mathcal{O}_{o}$ the ring of germs of holomorphic functions at $o$, and by $\mathfrak{m}$ the maximal ideal of $\mathcal{O}_{o}$. As mentioned above, $\mathcal{O}_{o}$ is a Noetherian regular excellent domain of characteristic zero. For the scheme $X=\operatorname{Spec}\mathcal{O}_{o}$, the valuations on $X$ can be regarded as the valuations on the ring $\mathcal{O}_{o}$. Especially, the following theorem shows that the algebraic Zhou valuations on $\mathcal{O}_{o}$ and the analytic Zhou valuations defined on $\mathcal{O}_{o}$ in [4] (see also [5] for a characterization) basically coincide.
Theorem 11.2.
Let $\mathfrak{q}\subseteq\mathcal{O}_{o}$ be a nonzero ideal. Then a valuation $v$ on $\mathcal{O}_{o}$ is an algebraic Zhou valuation related to $\mathfrak{q}$ with $v(\mathfrak{m})>0$ if and only if $v$ is an analytic Zhou valuation related to $|\mathfrak{q}|^{2}$.
The algebraic Zhou valuations related to $\mathfrak{q}$ are defined by Definition 3.2, and the definition of analytic Zhou valuations will be recalled in the following.
11.1.1. Zhou weights and Zhou valuations
We first recall the Zhou weights and Zhou valuations introduced in [4] (with some slight reformulations; see also [5]).
Definition 11.3 ([4, Definition 1.2, and Definition 1.18]).
Let $\mathfrak{q}\subseteq\mathcal{O}_{o}$ be a nonzero ideal. A psh germ $\Phi$ is called a Zhou weight related to $|\mathfrak{q}|^{2}$, if the following statements hold:
  1. (1)
    There exists a constant $N>0$ such that $\Phi\geq N\log|z|+O(1)$ near $o$;
  2. (2)
    $|\mathfrak{q}|^{2}e^{-2\Phi}$ is not integrable near $o$;
  3. (3)
    For any psh germ $\Psi$ satisfying that $|\mathfrak{q}|^{2}e^{-2\Psi}$ is not integrable near $o$ and $\Psi\geq\Phi+O(1)$ near $o$, we must have $\Psi=\Phi+O(1)$ near $o$.
The cone of Zhou weights consists of all psh germs $\Phi^{\prime}$ at $o$ such that there exists some constant $c>0$ and a Zhou weight $\Phi$ related to $|\mathfrak{q}|^{2}$ for some nonzero ideal $\mathfrak{q}\subseteq\mathcal{O}_{o}$, such that $\Phi^{\prime}=c\Phi+O(1)$ near $o$. In addition, one can also replace a Zhou weight by a maximal psh germ with the same singular type, i.e., there exists a psh germ $\widetilde{\Phi}=\Phi+O(1)$, such that $(\operatorname{dd^{c}}\widetilde{\Phi})^{n}\equiv 0$ on $U\setminus\{o\}$, where $U$ is an open neighborhood of $o$; see [4, Proposition 1.12]. For a maximal weight $\varphi$ near $o$, the relative type introduced by Rashkovskii in [41] is defined as:
$\sigma(u,\varphi)\coloneqq\sup\{c\geq 0\colon u\leq c\varphi+O(1)\text{ near }o\},$
where $u$ is any psh germ at $o$, and the maximality of $\varphi$ indicates (see [41, Proposition 3.4])
$u\leq\sigma(u,\varphi)\varphi+O(1)\quad\text{near }o.$
In particular, the relative types of Zhou weights are called Zhou numbers in [4].
Now let $v$ be a valuation on the ring $\mathcal{O}_{o}$.
Definition 11.4 ([4]).
We call $v$ an analytic Zhou valuation related to $|\mathfrak{q}|^{2}$, where $\mathfrak{q}\subseteq\mathcal{O}_{o}$ is a nonzero ideal, if there exists a Zhou weight $\Phi$ related to $\mathfrak{q}$ such that
$v(f)=\sigma(\log|f|,\Phi),\quad\forall\,(f,o)\in\mathcal{O}_{o}.$
Non-trivially, it was shown in [4, Corollary 1.9] that every Zhou weight can be associated with a Zhou valuation on $\mathcal{O}_{o}$ by establishing an integral expression form of Zhou numbers, where the proof involves a lot of analysis techniques.
11.1.2. Algebraic and analytic Zhou valuations
Now we turn back to prove Theorem 11.2. Before presenting the proof, let us briefly compare the logical structure of our proof with the proofs in [4]. The authors of [4] first introduced the notion of Zhou weight. They then proved that the Zhou numbers (the relative types to the Zhou weight) of the psh germs $u=\log|f|$ satisfy the definition of valuations (see [4, Corollary 1.7]), which leads to the notion of Zhou valuations. However, in our article, since we directly define what are Zhou valuations in the valuation space (see Definition 1.3), we must prove that, for each (algebraic) Zhou valuation, there exists a psh weight, which should be a Zhou weight, such that the relative type to the weight is compatible with the (algebraic) Zhou valuation. Although the construction of the weight in our proof is very similar with the constructions in [4, 5], but the proofs in [4, 5] strongly depend on the fact that, Zhou weights admit Zhou valuations, which will not be assumed in our proof.
In addition, we will avoid using analytical methods as much as possible in the proof. For example, the strong openness property of multiplier ideal sheaves ([26]; see also [28, 37]) will only be used to deduce $|\mathfrak{q}|^{2}e^{-2\Phi}$ is not integrable near oo” from $c_{o}^{\mathfrak{q}}(\Phi)=1$”. Meanwhile, a lower semi-continuity property of jumping numbers (see [27, Proposition 1.8]; see also [28]) is needed, where a special case, that is the lower semi-continuity property of complex singularity exponents, was proved in [16, Main Theorem 0.2] by Demailly’s approximation and the Ohsawa–Takegoshi $L^{2}$ extension theorem.
We start from a general valuation on $\operatorname{Spec}\mathcal{O}_{o}$. We will give a maximal psh weight associated with the valuation and establish some properties of the weight.
Denote the unit ball centered at $o$ in $\mathbf{C}^{n}$ by $\mathbf{B}(o,1)$.
Lemma 11.5.
Let $v$ be a valuation on $\mathcal{O}_{o}$ with $v(\mathfrak{m})>0$ and $A(v)<+\infty$. For every $z\in\mathbf{B}(o,1)$, set
$\Phi_{v}(z)\coloneqq\sup\Big\{\frac{\log|f(z)|}{v(f)}\colon f\in\mathcal{O}\big(\mathbf{B}(o,1)\big),\,f(o)=0,\,f\not\equiv 0,\,\sup_{\mathbf{B}(o,1)}|f|\leq 1\Big\}.$ (11.2)
Then the upper semi-continuous regularization $\Phi^{\star}_{v}$ of $\Phi_{v}$ is a negative psh function on $\mathbf{B}(o,1)$ satisfying the following statements:
  1. (1)
    There exist $C_{1},C_{2}>0$ such that $C_{1}\log|z|\leq\Phi_{v}^{\star}\leq C_{2}\log|z|$ near $o$.
  2. (2)
    The psh function $\Phi_{v}^{\star}$ is locally bounded and maximal on $\mathbf{B}(o,1)\setminus\{o\}$, and $e^{\Phi_{v}^{\star}}$ is continuous on $\mathbf{B}(o,1)$.
  3. (3)
    If $\mathfrak{q},\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}$ are nonzero ideals, and $\lambda\in(-\varepsilon_{0},0]\subseteq\mathbf{R}$ with $\varepsilon_{0}>0$ sufficiently close to 0, then
    $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}^{v}_{\bullet})=c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star}).$
    Especially, $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=c_{o}^{\mathfrak{q}}(\Phi_{v}^{\star})$ if $\lambda=0$.
  4. (4)
    If $\mathfrak{q},\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}$ are nonzero ideals, then the concave function (see Lemma 6.1)
    $\mathcal{T}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}^{v}_{\bullet}),\quad t\in(-\varepsilon_{0},+\infty)$
    satisfies that $\mathcal{T}^{\prime}_{-}(0)\geq\sigma(\log|\mathfrak{q}^{\prime}|,\Phi_{v}^{\star})$, where $\mathcal{T}^{\prime}_{-}(0)$ denotes the left derivative of $\mathcal{T}$ at 0, and $\sigma(\,\cdot\,,\Phi_{v}^{\star})$ is the relative type to $\Phi_{v}^{\star}$.
Proof.
Clearly $\Phi_{v}^{\star}$ is a negative psh function on $\mathbf{B}(o,1)$, and $\Phi_{v}^{\star}$ a.e. equals to $\Phi_{v}$ (cf. [11, Proposition 4.24]). Denote
$\mathcal{H}\coloneqq\Big\{f\in\mathcal{O}\big(\mathbf{B}(o,1)\big)\colon f(o)=0,\,f\not\equiv 0,\,\sup_{\mathbf{B}(o,1)}|f|\leq 1\Big\}.$
(1) Since $A(v)<+\infty$ and $v(\mathfrak{m})>0$, according to the Izumi’s inequality (ref. [30] or [31, Proposition 5.10]), we have
$v(\mathfrak{m})\operatorname{ord}_{o}(f)\leq v(f)\leq A(v)\operatorname{ord}_{o}(f)$
for every $f\in\mathcal{H}$. On the one hand, we have
$\Phi^{\star}_{v}(z)\geq\Phi_{v}(z)\geq\max_{1\leq i\leq n}\frac{\log|z_{i}|}{v(z_{i})}\geq\frac{1}{v(\mathfrak{m})}\max_{1\leq i\leq n}\log|z_{i}|\geq C_{1}\log|z|$
for some constant $C_{1}>0$, where $z=(z_{1},\cdots,z_{n})$. On the other hand, for every $f\in\mathcal{H}$,
$\frac{\log|f(z)|}{v(f)}\leq\frac{\operatorname{ord}_{o}(f)\log|z|}{v(f)}\leq\frac{\operatorname{ord}_{o}(f)\log|z|}{A(v)\operatorname{ord}_{o}(f)}=\frac{1}{A(v)}\log|z|,$
which implies $\Phi_{v}\leq C_{2}\log|z|$ with $C_{2}\coloneqq 1/A(v)$ and thus $\Phi_{v}^{\star}\leq C_{2}\log|z|$. The first statement holds.
(2) It follows from (1) that $\Phi_{v}^{\star}$ is locally bounded outside $o$ and $e^{\Phi_{v}^{\star}}$ is continuous. The standard arguments give the maximality of $\Phi_{v}^{\star}$.
(3) Since $A(v)<+\infty$, we have $\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})<+\infty$. By the coherence of ideal sheaves on $\mathbf{C}^{n}$, the germs of functions in $\mathcal{H}$ at $o$ generate every nonzero ideal of $\mathcal{O}_{o}$. Thus, we have
$\Phi_{v}^{\star}\geq\frac{1}{m}\log|\mathfrak{a}^{v}_{m}|+O(1)\quad\text{near }o,$ (11.3)
for every $m\geq 1$. Since
$m\cdot\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{m}^{v})=m\cdot c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}\left(\log|\mathfrak{a}_{m}^{v}|\right)=c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}\left(\frac{1}{m}\log|\mathfrak{a}_{m}^{v}|\right)$ (11.4)
by Proposition 11.1(3), letting $m\to+\infty$, it follows from (11.3) that
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})\leq c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star}).$
On the other hand, Choquet’s Lemma (cf. [11, Proposition 4.23]) shows that there exists a sequence $\phi_{j}=\log|f_{j}|/v(f_{j})$ with $f_{j}\in\mathcal{H}$ such that the $\Phi_{v}^{\star}=(\sup_{j}\phi_{j})^{\star}$. Since $f^{k}\in\mathcal{H}$ when $f\in\mathcal{H}$ for $k\in\mathbf{Z}_{+}$ and $v(f^{k})=kv(f)$, after replacing $f_{j}$ with some power of $f_{j}$, we may assume that
  1. (1)
    $v(f_{j})>1$ for all $j\geq 1$;
  2. (2)
    $v(f_{j+1})>v(f_{j})$ for all $j$ and $v(f_{j})\to+\infty$ when $j\to+\infty$;
  3. (3)
    $v(f_{j})/v(f_{j-1})\to+\infty$ when $j\to+\infty$.
Let $\varphi_{j}(z)\coloneqq\max\limits_{1\leq k\leq j}\phi_{k}(z)$. Then $\varphi_{j}$ increasingly a.e. converges to $\Phi_{v}^{\star}$, which implies $\sup\limits_{j\geq 1}c_{o}^{\mathfrak{q},\lambda\mathfrak{q}^{\prime}}(\varphi_{j})=c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star})$ when $\lambda\leq 0$, due to the lower semi-continuity property of jumping numbers (see [27, Proposition 1.8]).
Denote
$m_{j}=\lceil v(f_{j})+v(f_{j-1})\rceil,\quad m^{\prime}_{j}=\lceil v(f_{j})\rceil-1,\quad\forall\,j\geq 1,$
and
$p_{k,j}=\left\lceil\frac{m^{\prime}_{j}}{v(f_{k})}\right\rceil,\quad k=1,\ldots,j,$
which is a positive integer such that $g_{k,j}\coloneqq f_{k}^{p_{k,j}}\in\mathfrak{a}_{m^{\prime}_{j}}^{v}$ for each $k\leq j$. Observe that when $1\leq k\leq j-1$,
$v(g_{k,j})=p_{k,j}v(f_{k})$ $\leq\left(\frac{m^{\prime}_{j}}{v(f_{k})}+1\right)v(f_{k})$
$=m^{\prime}_{j}+v(f_{k})\leq v(f_{j})+v(f_{j-1})\leq m_{j},$
and $v(g_{j,j})=v(f_{j})\leq m_{j}$. It follows that
$\varphi_{j}$ $=\max_{1\leq k\leq j}\frac{\log|f_{k}|}{v(f_{k})}=\max_{1\leq k\leq j}\frac{\log|g_{k,j}|}{v(g_{k,j})}$
$\leq\frac{1}{m_{j}}\max_{1\leq k\leq j}\log|g_{k,j}|\leq\frac{1}{m_{j}}\log|\mathfrak{a}_{m_{j}^{\prime}}^{v}|+O(1)$
$=\frac{m^{\prime}_{j}}{m_{j}}\cdot\left(\frac{1}{m^{\prime}_{j}}\log|\mathfrak{a}_{m^{\prime}_{j}}^{v}|+O(1)\right)\quad\text{near }o,$
for every $j\geq 1$. Hence, since $\frac{m^{\prime}_{j}}{m_{j}}\to 1$ as $j\to+\infty$, according to (11.4), we obtain when $\lambda\in(-\varepsilon_{0},0]$,
$\mathrm{lct}(\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})=\sup_{j\geq 1}c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}\left(\frac{1}{m^{\prime}_{j}}\log|\mathfrak{a}_{m^{\prime}_{j}}^{v}|\right)\geq\sup_{j\geq 1}c_{o}^{\mathfrak{q},\lambda\mathfrak{q}^{\prime}}(\varphi_{j})=c_{o}^{\mathfrak{q},\lambda\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star}).$
Eventually, we get that the third statement holds.
(4) By Statement (1), the relative type
$\sigma\coloneqq\sigma(\log|\mathfrak{q}^{\prime}|,\Phi_{v}^{\star})\in(0,+\infty).$
Since $\Phi_{v}^{\star}$ is maximal, we have
$\log|\mathfrak{q}^{\prime}|\leq\sigma\Phi_{v}^{\star}+O(1)\quad\text{near }o.$
For $\mu>0$ sufficiently small and any $c>0$, it follows that
$\int_{V}|\mathfrak{q}|^{2}|\mathfrak{q}^{\prime}|^{-2\mu}e^{-2c\Phi_{v}^{\star}}=\int_{V}|\mathfrak{q}|^{2}e^{-2\mu\log|\mathfrak{q}^{\prime}|-2c\Phi_{v}^{\star}}\geq C_{0}\int_{V}|\mathfrak{q}|^{2}e^{-2(c+\mu\sigma)\Phi_{v}^{\star}},$
for any neighborhood $V$ of $o$ and $C_{0}>0$. Thus,
$c_{o}^{\mathfrak{q}}(\Phi_{v}^{\star})-\mu\sigma\geq c_{o}^{\mathfrak{q},-\mu\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star}),$
which implies
$\mathcal{T}(-\mu)=\mathrm{lct}(\mathfrak{q},-\mu\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})=c_{o}^{\mathfrak{q},-\mu\cdot\mathfrak{q}^{\prime}}(\Phi_{v}^{\star})\leq c_{o}^{\mathfrak{q}}(\Phi_{v}^{\star})-\mu\sigma=\mathcal{T}(0)-\mu\sigma$
by applying Statement (3). According to the concavity and monotonicity of the function $\mathcal{T}(t)$, we can verify that $\mathcal{T}^{\prime}_{-}(0)$ exists and $\mathcal{T}^{\prime}_{-}(0)\geq\sigma$.
The proof is complete. ∎
For Zhou valuations on $\mathcal{O}_{o}$, we can prove that the function $\Phi_{v}^{\star}$ satisfies more properties as demonstrated in the following proposition.
Proposition 11.6.
Let $\mathfrak{q}\subseteq\mathcal{O}_{o}$ be a nonzero ideal, and $v$ be a valuation on $\operatorname{Spec}\mathcal{O}_{o}$. If $v$ is an algebraic Zhou valuation related to $\mathfrak{q}$ with $v(\mathfrak{m})>0$, then the maximal psh function $\Phi^{\star}_{v}$ associated to $v$ defined as which in Lemma 11.5 on the unit ball $\mathbf{B}(o,1)$ additionally satisfies the following statements:
  1. (1)
    The jumping number $c_{o}^{\mathfrak{q}}(\Phi_{v}^{\star})=1$.
  2. (2)
    The relative type of $\Phi_{v}^{\star}$ is compatible with $v$, i.e.,
    $\sigma(\log|\mathfrak{q}^{\prime}|,\Phi_{v}^{\star})=v(\mathfrak{q}^{\prime}),\quad\text{for any ideal }\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}.$
  3. (3)
    For every psh germ $\psi$ at $o$ with $c_{o}^{\mathfrak{q}}(\psi)=1$ and $\psi\geq\Phi_{v}^{\star}+O(1)$ near $o$, it must hold that $\psi=\Phi_{v}^{\star}+O(1)$ near $o$.
Proof.
The statement (3) in Lemma 11.5 and $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$ given by Proposition 3.5 indicate that $c_{o}^{\mathfrak{q}}(\Phi_{v}^{\star})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{v})=1$.
For the function $\mathcal{T}(t)=\mathrm{lct}(\mathfrak{q},t\cdot\mathfrak{q}^{\prime};\mathfrak{a}_{\bullet}^{v})$:
  1. (a)
    The Statement (4) in Lemma 11.5 shows $\mathcal{T}^{\prime}_{-}(0)\geq\sigma(\log|\mathfrak{q}^{\prime},\Phi_{v}^{\star})$;
  2. (b)
    Since $v$ is a Zhou valuation related to $\mathfrak{q}$, Theorem 8.1 implies $\mathcal{T}^{\prime}_{-}(0)=\mathcal{T}^{\prime}_{+}(0)=v(\mathfrak{q}^{\prime})$;
  3. (c)
    The definition of $\Phi_{v}^{\star}$ gives $\sigma(\log|\mathfrak{q}^{\prime}|,\Phi_{v}^{\star})\geq v(\mathfrak{q}^{\prime})$.
Thus, $\sigma(\log|\mathfrak{q}^{\prime}|,\Phi_{v}^{\star})=v(\mathfrak{q}^{\prime})$ for every nonzero ideal $\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}$.
Now we prove the third statement. Suppose that the psh germ $\psi$ at $o$ satisfies $c_{o}^{\mathfrak{q}}(\psi)=1$ and $\psi\geq\Phi_{v}^{\star}+O(1)$ near $o$. Denote
$\mathfrak{b}_{t}=\mathcal{J}(t\psi)_{o},\quad t\in(0,+\infty).$
Note that $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{t}\}_{t}$ is a subadditive system of ideals in $\mathcal{O}_{o}$ (cf. [15]). According to Proposition 11.1, we have $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})=c_{o}^{\mathfrak{q}}(\psi)=1$, and $\mathfrak{b}_{\bullet}$ has controlled growth.
Since $\psi\geq\Phi_{v}^{\star}+O(1)$ near $o$, we can verify that $\mathfrak{b}_{j}\supseteq\mathfrak{a}^{v}_{j}$ for every $j\in\mathbf{Z}_{+}$. In fact, by (11.3), for every $N\in\mathbf{Z}_{+}$,
$\psi\geq\frac{1}{N}\log|\mathfrak{a}_{N}^{v}|+O(1)\quad\text{near }o,$
which implies
$\mathfrak{b}_{j}=\mathcal{J}(j\psi)_{o}\supseteq\mathcal{J}\left(\frac{j}{N}\log|\mathfrak{a}_{N}^{v}|\right)_{o}=\mathcal{J}\left(\frac{j}{N}\cdot\mathfrak{a}_{N}^{v}\right)=\mathcal{J}(j\cdot\mathfrak{a}_{\bullet}^{v})$
by Proposition 11.1(3), where $N\gg 1$ is divisible enough. Especially, we have $\mathfrak{a}_{j}^{v}\subseteq\mathcal{J}(j\cdot\mathfrak{a}_{\bullet}^{v})\subseteq\mathfrak{b}_{j}$.
Since $v(\mathfrak{m})>0$ and $\mathfrak{a}_{j}^{v}\subseteq\mathfrak{b}_{j}$ for every $j$, there exists some positive integer $p$ such that $\mathfrak{m}^{pj}\subseteq\mathfrak{b}_{j}$ for every $j\in\mathbf{Z}_{+}$. Now according to [32, Proposition 2.1], we can find a valuation $w$ on $\operatorname{Spec}\mathcal{O}_{o}$ with $A(w)<+\infty$ which computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})$. After rescaling, we can assume that $w(\mathfrak{b}_{\bullet})=1$. Then it follows from Lemma 2.10 that
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a}_{\bullet}^{w})\leq A(w)+w(\mathfrak{q})=\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})=1.$
Since $\mathfrak{a}_{j}^{v}\subseteq\mathfrak{b}_{j}$ for every $j$, we have $w(\mathfrak{a}_{\bullet}^{v})\geq w(\mathfrak{b}_{\bullet})=1$, which implies $w\geq v$ by Lemma 2.5. Now, according to the assumption that $v$ is an algebraic Zhou valuation related to $\mathfrak{q}$, we get $w=v$, which indicates that $v(\mathfrak{b}_{\bullet})=w(\mathfrak{b}_{\bullet})=1$. Then it follows from [31, Proposition 6.5] and Proposition 11.1(2) that
$\frac{v(\mathfrak{b}_{t})}{t}\geq v(\mathfrak{b}_{\bullet})-\frac{A(v)}{t}\geq 1-\frac{1}{t},\quad\forall t>0,$
where $A(v)=1-v(\mathfrak{q})\leq 1$ by Theorem 4.1. It implies that
$\sigma(\log|\mathfrak{b}_{j}|,\Phi_{v}^{\star})=v(\mathfrak{b}_{j})\geq j-1$
when $j\in\mathbf{Z}_{\geq 2}$ by the statement (2), i.e.,
$\frac{1}{j}\log\big|\mathcal{J}(j\psi)_{o}\big|=\frac{1}{j}\log|\mathfrak{b}_{j}|\leq\frac{j-1}{j}\Phi_{v}^{\star}+O(1)\quad\text{near }o,$
for every $j\gg 1$. Finally, we apply Demailly’s approximation theorem (or the Ohsawa-Takegoshi $L^{2}$ extension theorem) to $\psi$ to get
$\psi\leq\frac{1}{j}\log\big|\mathcal{J}(j\psi)_{o}\big|+O(1)\leq\frac{j-1}{j}\Phi_{v}^{\star}+O(1)\quad\text{near }o,$
which implies
$\sigma(\psi,\Phi_{v}^{\star})\geq\frac{j-1}{j}\to 1,\quad j\to+\infty.$
Hence, $\psi\leq\Phi_{v}^{\star}+O(1)$ near $o$, and thus $\psi=\Phi_{v}^{\star}+O(1)$ near $o$.
The proof is complete. ∎
Now we prove Theorem 11.2.
Proof of Theorem 11.2.
We first assume that $v$ is an algebraic Zhou valuation on $\operatorname{Spec}\mathcal{O}_{o}$ with $v(\mathfrak{m})>0$. Then Lemma 11.5 and Proposition 11.6 show that there exists a maximal psh germ $\Phi$ at $o$ such that the following statements hold:
  1. (a)
    There exist positive constants $C_{1},C_{2}$ such that $C_{1}\log|z|\leq\Phi\leq C_{2}\log|z|$ near $o$.
  2. (b)
    $c_{o}^{\mathfrak{q}}(\Phi)=1$.
  3. (c)
    For every psh germ $\psi$ at $o$ with $c_{o}^{\mathfrak{q}}(\psi)=1$ and $\psi\geq\Phi+O(1)$ near $o$, it must be $\psi=\Phi+O(1)$ near $o$.
  4. (d)
    $\sigma(\log|\mathfrak{q}^{\prime}|,\Phi)=v(\mathfrak{q}^{\prime})$ for every ideal $\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}$.
The statement (b) and the strong openness property ([26]) show that $|\mathfrak{q}|^{2}e^{-2\Phi}$ is not integrable near $o$. Now we can see that the statements (a) (b) and (c) actually tell that $\Phi$ is a Zhou weight related to $|\mathfrak{q}|^{2}$ near $o$; see Definition 11.3. Then it follows from the statement (d) that, restricted to $\mathcal{O}_{o}$, $v$ is an analytic Zhou valuation related to $|\mathfrak{q}|^{2}$ by Definition 11.4.
Next, we prove the converse side. In fact, the proof is quite similar with the proof of the statement (3) of Proposition 11.6.
Let $v$ be a valuation on $\operatorname{Spec}\mathcal{O}_{o}$, such that $v$ is an analytic Zhou valuation related to $|\mathfrak{q}|^{2}$. By Definition 11.3 and Definition 11.4 again, we assume that $\Psi$ is a local Zhou weight (and maximal) associated to the valuation $v$, i.e.,
$\sigma(\log|f|,\Psi)=v(f),\quad\forall(f,o)\in\mathcal{O}_{o}.$ (11.5)
Set
$\mathfrak{b}_{t}=\mathcal{J}(t\Psi)_{o},\quad t\in(0,+\infty).$
[…]
Let $w\geq\tilde{v}$ be an algebraic Zhou valuation on $\operatorname{Spec}\mathcal{O}_{o}$ related to $\mathfrak{q}$, and then it suffices to show $w=v$. Define the function $\Phi_{w}^{\star}$ associated with $w$ by (11.2), where
$\Phi_{w}(z)\coloneqq\sup\Big\{\frac{\log|f(z)|}{w(f)}\colon f\in\mathcal{O}\big(\mathbf{B}(o,1)\big),\,f(o)=0,\,f\not\equiv 0,\,\sup_{\mathbf{B}(o,1)}|f|\leq 1\Big\}.$
Then $c_{o}^{\mathfrak{q}}(\Phi_{w}^{\star})=1$ by Proposition 11.6(1). According to Proposition 11.6(2) and (11.5), to prove $w=v$, we only need to prove $\Psi=\Phi_{w}^{\star}+O(1)$ near $o$. Because $\Psi$ is a Zhou weight related to $|\mathfrak{q}|^{2}$, it suffices to show $\Phi_{w}^{\star}\geq\Psi+O(1)$ near $o$.
Since $w\geq\tilde{v}$ and $\tilde{v}(\mathfrak{b}_{\bullet})=1$, we have $w(\mathfrak{b}_{\bullet})\geq 1$. Then it follows from [31, Proposition 6.5] (or Lemma 2.7 (1)) and Proposition 11.1(2) that
$\frac{w(\mathfrak{b}_{t})}{t}\geq w(\mathfrak{b}_{\bullet})-\frac{A(w)}{t}\geq 1-\frac{1}{t},\quad\forall t>0,$
as $A(w)=1-w(\mathfrak{q})\leq 1$. Using Proposition 11.6 and Demailly’s approximation theorem, we get
$\Phi^{\star}_{w}\geq\frac{1}{w(\mathfrak{b}_{j})}\log|\mathfrak{b}_{j}|+O(1)\geq\frac{j}{j-1}\Psi+O(1)\quad\text{near }o,$
for every $j>1$. This implies that the relative type
$\sigma(\Psi,\Phi_{w}^{\star})\geq 1-\frac{1}{j}\to 1\quad\text{as }j\to+\infty.$
Consequently, we get $\Psi\leq\Phi_{w}^{\star}+O(1)$ near $o$, and the proof is complete. ∎
We explain in the following corollaries and examples that how we can prove some other results in [4] which are not mentioned above by algebraic methods.
Corollary 11.7 is a very nontrivial result of the Zhou weights, established in [4] by giving an integral expression for the Zhou numbers (cf. [4, Theorem 1.6]), which highly relies on the analytic methods, and we can reprove it using the same arguments in the proofs of Proposition 11.6 and Theorem 11.2.
Corollary 11.7 ([4, Corollary 1.9]).
If $\Phi$ is a Zhou weight related to some nonzero ideal $\mathfrak{q}\subseteq\mathcal{O}_{o}$, then there exists a valuation $v$ on $\mathcal{O}_{o}$ such that
[…]
Proof.
Let $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{t}\}_{t>0}=\{\mathcal{J}(t\Phi)_{o}\}_{t>0}$ be the subadditive system associated with the Zhou weight $\Phi$. Note that there exists $N>0$ such that $\Phi\geq N\log|z|+O(1)$ near $o$. Then using [32, Proposition 2.1], we can find a valuation $w$ on $\mathcal{O}_{o}$ which computes $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{b}_{\bullet})$, and $w(\mathfrak{b}_{\bullet})=1$. Let $v\geq w$ be a Zhou valuation on $\mathcal{O}_{o}$ related to $\mathfrak{q}$, and $\Phi_{v}^{\star}$ be the Zhou weight associated with $v$ related to $|\mathfrak{q}|^{2}$. With the same arguments used in the proofs of Proposition 11.6 (3) and also Theorem 11.2, we can get $\Phi_{v}^{\star}\geq\Phi+O(1)$, and thus $\Phi_{v}^{\star}=\Phi+O(1)$ since both $\Phi_{v}^{\star}$ and $\Phi$ are Zhou weights related to $|\mathfrak{q}|^{2}$. Therefore, we have that $v(\mathfrak{q}^{\prime})=\sigma(\log|\mathfrak{q}^{\prime}|,\Phi)$ for every nonzero ideal $\mathfrak{q}^{\prime}\subseteq\mathcal{O}_{o}$. ∎
Corollary 11.8.
All Zhou weights are tame.
Proof.
This is a consequence of [4, Theorem 1.10], which was proved by analytic approach there. We give an algebraic proof in the following.
Recall that a psh weight $\Phi$ is said to be tame, if there exists a constant $C>0$ such that for every $t>0$ and every $(f,o)\in\mathcal{I}(t\varphi)_{o}$, $\log|f|\leq(t-C)\Phi+O(1)$ holds near $o$. In fact, observing that for a Zhou valuation $v$ related to a nonzero ideal $\mathfrak{q}$ and the associated Zhou weight $\Phi_{v}^{\star}$,
$c_{o}^{f}(\Phi_{v}^{\star})$ $=\mathrm{lct}^{(f)}(\mathfrak{a}_{\bullet}^{v})\leq\frac{A(v)+v(f)}{v(\mathfrak{a}_{\bullet}^{v})}=1-v(\mathfrak{q})+v(f)$
$=1-v(\mathfrak{q})+\sigma(\log|f|,\Phi_{v}^{\star}),$
we can thus directly conclude that the Zhou weights are tame using Theorem 11.2, Corollary 11.7 and the fact that $\Phi_{v}^{\star}$ is maximal. ∎
Example 11.9 (cf. [4, Example 1.5 & 1.20]).
Let $(k_{1},\ldots,k_{n})\in\mathbf{Z}_{+}^{n}$, and $\Phi(z)\coloneqq\max_{1\leq i\leq n}\frac{1}{\alpha_{i}}\log|z_{i}|$ be a psh function on $\Delta^{n}\subseteq\mathbf{C}^{n}$, where $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbf{R}_{>0}^{n}$ satisfying $\sum_{i=1}^{n}{\alpha_{i}}k_{i}=1$. Then $\Phi$ is a Zhou weight related to $|\mathfrak{q}|^{2}$ near $o$ for $\mathfrak{q}=(z_{1}^{k_{1}-1}\cdots z_{n}^{k_{n}-1})$.
Proof.
The proof of a special case of this example in [4] relies on the concavity property of minimal $L^{2}$ integrals (cf. [23, 24, 25]), which somehow is a deep analytic result. We give a mainly algebraic proof here.
Let $v_{\alpha}$ be the monomial valuation defined on $\mathcal{O}_{o}$ by
$v_{\alpha}(f)\coloneqq\min\{\langle\alpha,\beta\rangle\colon c_{\beta}\neq 0\},\quad\forall\,f(z)=\sum\limits_{\beta\in\mathbf{Z}^{n}_{\geq 0}}c_{\beta}z^{\beta}\in\mathcal{O}_{\mathbf{C}^{n},o}.$
Then $v_{\alpha}$ is a quasi-monomial adapted to $(X,H)=(\operatorname{Spec}\mathcal{O}_{o},H_{1}+\cdots+H_{n})$ with $H_{i}=V(z_{i})$, and $A(v_{\alpha})=\sum_{i=1}^{n}\alpha_{i}$ (cf. [32, Section 8]). Set
$\mathfrak{a}\coloneqq(z_{1}^{k_{1}}\cdots z_{n}^{k_{n}}).$
We have $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})=1$. In fact, on the one hand, we have
$\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\leq\frac{A(v_{\alpha})+v_{\alpha}(\mathfrak{q})}{v_{\alpha}(\mathfrak{a})}=\frac{\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\alpha_{i}(k_{i}-1)}{\sum_{i=1}^{n}\alpha_{i}k_{i}}=1.$
On the other hand, we have
$\mathrm{lct}(\mathfrak{a})=\inf_{w\in\mathrm{Val}_{X}^{*}}\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a})}$
by Lemma 2.10. Consider the retraction map $r_{X,H}\colon\mathrm{Val}_{X}\to\mathrm{QM}(X,H)$. Then $\bar{w}\coloneqq r_{X,H}(w)$ is a monomial valuation. We have $w(\mathfrak{q})=\bar{w}(\mathfrak{q})$, $w(\mathfrak{a})=\bar{w}(\mathfrak{a})$ and $A(w)\geq A(\bar{w})$, which implies
$\frac{A(w)+w(\mathfrak{q})}{w(\mathfrak{a})}\geq\frac{A(\bar{w})+\bar{w}(\mathfrak{q})}{\bar{w}(\mathfrak{a})}=\frac{\sum_{i=1}^{n}\bar{w}_{i}+\sum_{i=1}^{n}\bar{w}_{i}(k_{i}-1)}{\sum_{i=1}^{n}\bar{w}_{i}k_{i}}=1,$
and thus $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})\geq 1$, where $\bar{w}_{i}=\bar{w}(z_{i})$.
Now we can conclude that $\mathrm{lct}^{\mathfrak{q}}(\mathfrak{a})$ is computed by the valuation $v_{\alpha}$, and $v_{\alpha}(\mathfrak{a})=\sum_{i=1}^{n}\alpha_{i}k_{i}=1$. Therefore, by Example 4.4, $v_{\alpha}$ is a Zhou valuation related to $\mathfrak{q}$, and it is easy to check that $\Phi$ is a psh weight whose relative type (which actually is the Kiselman number; see [33]) is compatible with $v_{\alpha}$. Hence, $\Phi$ is a Zhou weight related to $|\mathfrak{q}|^{2}$ near $o$. ∎
There are also several results obtained in [5] based on the notion of $\mathscr{A}$” introduced there. According to Theorem 4.1 and Theorem 11.2, it is easy to see that the notion $\mathscr{A}(v)$” defined in [5] in fact coincides with $A(v)$, the log-discrepancy, for all (analytic) Zhou valuations $v$ on $\mathcal{O}_{o}$. Consequently, we can replace all the $\mathscr{A}$” by “AA” in [5], which will provide some corollaries, but we do not demonstrate for detail in the present paper.
Acknowledgments
The authors would like to thank Mattias Jonsson, Zhitong Mi, Xun Sun, Zheng Yuan and Xiangyu Zhou for helpful discussions. The second named author was supported by National Key R&\&D Program of China 2021YFA1003100, NSFC-11825101 and NSFC-12425101. The third named author was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 948066 (ERC-StG RationAlgic).
Appendix A The two-dimensional case
After an early preprint version of the present paper was released, Jonsson asked the authors the following question in a private communication:
Question A.1.
Can one prove that every quasi-monomial valuation is a Zhou valuation related to some ideal $\mathfrak{q}$?
Until now, we have not been able to answer this question in general. However, by using the valuative tree theory developed by Favre and Jonsson (cf. [20, 21, 22]), we can show that every nontrivial quasi-monomial valuation on a $2$-dim regular excellent scheme $X$ over $\mathbf{Q}$ is a Zhou valuation related to some ideal $\mathfrak{q}$ (up to rescaling), i.e., $\mathrm{ZVal}_{X}=\mathrm{QM}_{X}\setminus\{\text{trivial valuation}\}$ for $\dim X\leq 2$ (see Corollary 4.7).
Let $v$ be a nontrivial quasi-monomial valuation on $X$. If $\xi=c_{X}(v)$ is not a closed point, then $\xi$ is a generic point of an irreducible curve. Since $\dim\mathcal{O}_{X,\xi}=1$, we can see $v=\lambda\cdot\operatorname{ord}_{\xi}$ for some constant $\lambda>0$. Of course, such valuations must be Zhou valuations by Example 4.4. Hence, we only need to consider the case that $\xi$ is a closed point. According to [31, Proposition 7.12 & Lemma 3.10] and Cohen’s structure theorem, we may assume $X=\operatorname{Spec}\widehat{\mathcal{O}_{X,\xi}}=\operatorname{Spec}k[[x,y]]$ for $\xi=(0,0)$ and a field $k$ of characteristic 0. Furthermore, by [31, Proposition 7.13 & Lemma 3.11], it suffices to consider the case that $k$ is algebraically closed. After rescaling, we can normalized $v$ by $v(\mathfrak{m})=1$, where $\mathfrak{m}=(x,y)$ the maximal ideal defining the closed point $(0,0)$.
Recall that the set consisting of such normalized centered quasi-monomial valuations on $R=k[[x,y]]$, denoted by $\mathcal{V}_{\mathrm{qm}}$, is a subtree of the complete tree $\mathcal{V}$ of normalized centered valuations (may take values in $\mathbf{R}_{+}\cup\{\infty\}$), where $\mathcal{V}$ is the so-called valuative tree (see [20]). Therefore, it suffices to prove the following proposition, and in the proof, we will adopt some notations and conventions which can be referred to [20, 21].
Proposition A.2.
For every quasi-monomial valuation $v\in\mathcal{V}_{\mathrm{qm}}$, there exists a large enough positive integer $N\gg 1$, such that $v$ is a Zhou valuation related to $\mathfrak{m}^{N}$ up to rescaling.
Proof.
Due to Theorem 4.1 and Corollary 4.2, (as explained by the paragraph after Corollary 4.2) we only need to prove that there exists $N\gg 1$ such that
$\big\{w\in\mathcal{V}\colon w\ \text{computes}\ \mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})\big\}=\{v\}.$ (A.1)
Fix $N\in\mathbf{Z}_{\geq 0}$, which will be determined at last. We assume $v\neq w\in\mathcal{V}$ and $w$ computes $\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})$, i.e.,
$\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})=\inf_{\widetilde{v}\in\mathcal{V}_{\mathrm{qm}}}\frac{A(\widetilde{v})+N}{\widetilde{v}(\mathfrak{a}_{\bullet}^{v})}=\frac{A(w)+N}{w(\mathfrak{a}_{\bullet}^{v})}.$
We denote the last term in the above equality by $\sigma(w)$. If $w>v$, then $A(w)>A(v)$ by [20, Section 3.6]. In addition, $w(\mathfrak{a}_{\bullet}^{v})=1$ since $w>v$ and $w(\mathfrak{m})=v(\mathfrak{m})=1$. It follows that $\sigma(w)>\sigma(v)\geq\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})$, which shows $w$ can not compute $\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})$.
Now we assume $w>v$ does not hold. Then the infimum $w\wedge v<v$, and either $w\wedge v=w$ or $w\wedge v<w$ holds. Due to [20, Proposition 3.25] and the tree structure of $\mathcal{V}$, we have
$w(\mathfrak{a}_{\bullet}^{v})=\inf_{\phi}\frac{w(\phi)}{v(\phi)}=\inf_{\phi}\frac{m(\phi)\alpha(w\wedge v_{\phi})}{m(\phi)\alpha(v\wedge v_{\phi})}=\inf_{\phi}\frac{\alpha(w\wedge v_{\phi})}{\alpha(v\wedge v_{\phi})}=\frac{\alpha(w\wedge v)}{\alpha(v)},$
where the infimum are over all irreducible $\phi\in\mathfrak{m}$, $m(\phi)$ is the multiplicity of $\phi$, $\alpha$ is the skewness (cf. [20, Definition 3.23]), and $v_{\phi}$ is the curve valuation (cf. [20, Section 1.5.5]) defined by the curve $C=\{\phi=0\}$. Therefore, we have
$\sigma(w)=\frac{A(w)+N}{\alpha(w\wedge v)}\alpha(v)>\frac{A(w\wedge v)+N}{\alpha(w\wedge v)}\alpha(v)=\sigma(w\wedge v)$
once $w>w\wedge v$ does not hold, which shows that this is also not the case.
According to the above discussions, we can assume $w<v$, that is, $w\in\mathcal{V}_{\mathrm{qm}}$ lying in the segment $v_{\mathfrak{m}},v)$, where the multiplicity valuation $v_{\mathfrak{m}}=m(\,\cdot\,)$ is the root of the tree. Let
$v_{\mathfrak{m}}=v_{0}<v_{1}<\ldots<v_{g}<v_{g+1}=v$
be the approximation sequence (cf. [20, Definition 3.45]) associated to $v$, which is finite since $v\in\mathcal{V}_{\mathrm{qm}}$. Then (cf. [20, Section 3.6])
$A(v)=2+\sum_{j=1}^{g+1}m_{j}(\alpha_{j}-\alpha_{j-1}),$
[…]
$A(t)=2+m_{g+1}(t-\alpha_{g})+\sum_{j=1}^{g}(\alpha_{j}-\alpha_{g-1})$
is a linear function in $t\in(\alpha_{g},\alpha_{g+1}]$, and hence,
$\frac{\mathrm{d}\sigma(t)}{\mathrm{d}t}=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{A(t)+N}{t}\cdot\alpha(v)\right)=\alpha(v)\frac{m_{g+1}\cdot t-A(t)-N}{t^{2}},$ (A.2)
for $t\in(\alpha_{g},\alpha_{g+1}]$. Thus, if we take
[…]
then $\sigma(t)$ is strictly decreasing in $t\in(\alpha_{g},\alpha_{g+1}]$, i.e., any $w\in(v_{g},v_{g+1})$ can not compute $\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})$. Continuously, we can see that if we take
[…]
then $\sigma(t)$ is strictly decreasing in $t\in[1,v_{g+1}]$, that is, the valuation $w$ computing $\mathrm{lct}^{\mathfrak{m}^{N}}(\mathfrak{a}_{\bullet}^{v})$ also can not lie in the segment $v_{\mathfrak{m}},v)$.
In conclusion, for every positive integer NN satisfying (A.3), we have verified that (A.1) holds, which implies that $v$ is a Zhou valuation related to the ideal $\mathfrak{m}^{N}$ up to rescaling. ∎
In particular, for the trivial ideal $\mathfrak{q}=(1)$ we have the following:
Proposition A.3.
The subset $\mathrm{ZV}(1)$ of $\mathcal{V}_{\mathrm{qm}}$ consisting of Zhou valuations related to the trivial ideal $\mathfrak{q}=(1)$ is given by:
$\mathrm{ZV}(1)=\big\{v\in\mathcal{V}_{\mathrm{qm}}\colon m(v)=1\big\}.$
Proof.
The argument used in the first part of the proof of Proposition A.2 also shows that
$\mathrm{ZV}(1)=\big\{v\in\mathcal{V}_{\mathrm{qm}}\colon\sigma_{0}(w)>\sigma_{0}(v)\ \text{for all}\ w\in\mathcal{V}_{\mathrm{qm}}\ \text{with}\ w<v\big\},$
where
$\sigma_{0}(w)\coloneqq\frac{A(w)}{w(\mathfrak{a}_{\bullet}^{v})}=\frac{A(w)}{\alpha(w)}\alpha(v),$
for $w<v$.
If $m(v)=1$, then for any $w<v$, we have $m(w)=1$, so
$\sigma_{0}(w)=\frac{1+\alpha(w)}{\alpha(w)}\alpha(v)>1+\alpha(v)=\sigma_{0}(v),$
by [20, Proposition 3.48]. Hence, any $v\in\mathcal{V}_{\mathrm{qm}}$ with $m(v)=1$ belongs to $\mathrm{ZV}(1)$.
Conversely, if $m(v)>1$, we can take
$v_{g}\coloneqq\max\big\{w\in[v_{\mathfrak{m}},v)\colon m(w)<m(v)\big\},$
which appears in the approximation sequence of $v$ satisfying $v_{g}<v$. Set $\alpha_{g}=\alpha(v_{g})<\alpha(v)=\alpha_{g+1}$. Due to (A.2), for $t\in(\alpha_{g},\alpha_{g+1}]$, we have
$\frac{\mathrm{d}\sigma_{0}(t)}{\mathrm{d}t}=\alpha(v)\frac{m(w_{t})\alpha(w)-A(w_{t})}{\alpha(w_{t})^{2}},$
where $t=\alpha(w_{t})$ for $w_{t}\in(v_{g},v]$. However, it follows from [20, Proposition 3.48] that $A(w_{t})<m(w_{t})\alpha(w_{t})$ as $m(w_{t})=m(v)>1$ for all $w_{t}\in(v_{g},v]$. Thus, the left derivative of $\sigma_{0}$ at $t=\alpha(v)$ is strictly positive, which shows that $v$ does not attain the infimum of $\sigma_{0}(\,\cdot\,)$ in $\mathcal{V}_{\mathrm{qm}}$ (in fact, one can show that the infimum is attained at $v_{1}=\max\{w\colon w\leq v,\ m(w)=1\}$). Therefore, $v\notin\mathrm{ZV}(1)$ when $m(v)>1$. ∎